### v1

# Post-reduction inference for confidence sets of models

Post-reduction inference for confidence sets of models
Heather Battey Heather Battey, Yanbo Tang: Department of Mathematics, Imperial College London, 180 Queen’s Gate, London, SW7 2AZ, UK h.battey@imperial.ac.uk, yanbo.tang@imperial.ac.uk Daniel García Rasines Daniel García Rasines: Colegio Universitario de Estudios Financieros, Calle Almansa 101, Madrid 28040, Spain. daniel.garciarasines@cunef.edu  and  Yanbo Tang
Summary.
Sparsity in a regression context makes the model itself an object of interest, pointing to a confidence set of models as the appropriate presentation of evidence. A difficulty in areas such as genomics, where the number of candidate variables is vast, arises from the need for preliminary reduction prior to the assessment of models. The present paper considers a resolution using inferential separations fundamental to the Fisherian approach to conditional inference, namely, the sufficiency/co-sufficiency separation, and the ancillary/co-ancillary separation. The advantage of these separations is that no direction for departure from any hypothesised model is needed, avoiding issues that would otherwise arise from using the same data for reduction and for model assessment. In idealised cases with no nuisance parameters, the separations extract all the information in the data, solely for the purpose for which it is useful, without loss or redundancy. The extent to which estimation of nuisance parameters affects the idealised information extraction is illustrated in detail for the normal-theory linear regression model, extending immediately to a log-normal accelerated-life model for time-to-event outcomes. This idealised analysis provides insight into when sample-splitting is likely to perform as well as, or better than, the co-sufficient or ancillary tests, and when it may be unreliable. The considerations involved in extending the detailed implementation to canonical exponential-family and more general regression models are briefly discussed. As part of the analysis for the Gaussian model, we introduce a modified version of the refitted cross-validation estimator of Fan et al. (2012), whose distribution theory is exact in an appropriate conditional sense.
Some key words: Ancillarity; Co-sufficient manifold; Inferential separations; Information geometry; Model uncertainty; Regression; Selective inference; Sparsity.
1. Introduction
For regression problems with many potential explanatory variables, Cox (1968) noted that multiple low-dimensional combinations of variables are often compatible with the data. Although an arbitrary choice between well-fitting models is reasonable for prediction, this will typically be misleading in scientific studies in which subject-matter understanding is sought. This perspective, which was emphasised repeatedly over many years (e.g. Cox, 1995; Cox and Snell, 1974, 1989), has acquired renewed importance from a multitude of modern problems, notably in genomics, in which the number of variables for consideration is vast. Routine practice is to report a single model identified by a variable-selection algorithm, the most widely used being the lasso (Tibshirani, 1996). The recommendation of Cox and Battey (2017), by constrast, is to report all models that are compatible with the data at a chosen significance level, i.e. a confidence set of models. Some theoretical considerations associated with this proposal under a high-dimensional regime were developed by Battey and Cox (2018) and Lewis and Battey (2025).
Suppose that a regression model in $p$ variables is correctly specified but over-conditioned, in the sense that the vector of regression coefficients contains many zero entries, whose positions are unknown. In the remainder of the paper, we therefore equate the specification of a model with a set of variable indices and write $\mathcal{E}^{*}\subset[p]$ for the sparse model in $s^{*}<p$ variables that generated the data. When $p$ is not so large as to make the procedure practically infeasible or theoretically unsound, a simple way of constructing confidence sets of models is via a likelihood-ratio test of every low-dimensional subset of variables $\mathcal{E}_{m}\subset[p]:=\{1,\ldots,p\}$ against the encompassing model $[p]$. All models $\mathcal{E}_{m}$, $m=1,\ldots,2^{p}$ not rejected at level $\alpha$ comprise the confidence set of models $\mathcal{M}$, and this, by standard arguments, contains $\mathcal{E}^{*}$ with asymptotic probability $1-\alpha$ in hypothetical repeated use. It is sensible in practice to restrict attention only to those models of reasonable size, substantially reducing the number of assessments to be made.
For the genomics applications we have in mind, $p$ is typically of the order of several thousand, while the number $n$ of individuals is usually in the hundreds or even dozens. The above construction of $\mathcal{M}$ is therefore infeasible without preliminary reduction from $[p]$ to an encompassing model, $\hat{\mathcal{E}}$ say, of more manageable size. The circumflex emphasises that $\hat{\mathcal{E}}$ is constructed from the data; it is not intended as an estimator of $\mathcal{E}^{*}$, a much smaller subset. Any reasonable variable screening procedure can be used for the construction of $\hat{\mathcal{E}}$; see Lewis and Battey (2025) for a theoretical discussion of the implications of that choice.
The key point as far as the present work is concerned is that, since the model $\hat{\mathcal{E}}$ is selected in the light of the data, it fits the observed data better than an arbitrary model of the same size encompassing the one to be tested. For this reason, a likelihood-ratio test of $\mathcal{E}_{m}\subset\hat{\mathcal{E}}$ against $\hat{\mathcal{E}}$ typically rejects too often in hypothetical repeated use, so that the resulting model confidence set $\mathcal{M}$ has a lower coverage probability than suggested by the nominal level of the test. Battey and Cox (2018) used sample splitting as a simple device for guaranteeing the nominal coverage, although it was noted that a potentially better approach would use a minimal sufficiency separation to assess model adequacy (Barndorff-Nielsen and Cox, 1994), bypassing the likelihood ratio test and obviating sample splitting. Essentially the same inferential separation was used by Engen and Lillegård (1997), Lindqvist et al. (2003), Lindqvist and Taraldsen (2005), Lockhart et al. (2007), and Barber and Janson (2022) to assess model fit, but not in the context of the post-reduction inference considered here, where the arguments in favour of inferential separations, relative to a likelihood-ratio or other directed test, are much stronger. The approaches taken in the above references are based on Monte Carlo algorithms to approximately sample from the appropriate subset of the sample space after suitable conditioning. Markov Chain Monte Carlo methods are too computationally expensive for present purposes, as construction of $\mathcal{M}$ entails assessing all low-dimensional subsets of $\hat{\mathcal{E}}$ for their compatibility with the data, a combinatorially large problem.
The principles developed in the present paper, specifically for the purpose of confidence sets of sparse regression models, are based on the sufficiency/co-sufficiency separation, and the ancillary/co-ancillary separation fundamental to the Fisherian approach to conditional inference and reviewed in Chapters 2.3 and 2.5 of Barndorff-Nielsen and Cox (1994). In the absence of a proposed direction of departure from the postulated model, a combination of the ancillary and co-sufficient components in principle achieves the maximal extraction of information relevant for the assessment of model adequacy, having extracted the information relevant for estimation and reduction. In practice there is considerable flexibility in how such information is used, the most naïve approaches being ineffectual for detecting relevant departures.
A key question is over the efficacy of sample splitting relative to the more sophisticated approaches developed here. In simpler contexts, methods powerful against specific alternatives are often valuable in small samples. This raises the possibility that using a large portion of the data for reduction and a small portion for model assessment may be as or more effective than the inferential separations considered here, provided that the test for model assessment specifies a direction for departure from the null, and provided that the sample size for model assessment is sufficiently large that any asymptotic distributions hold to an adequate order of approximation. In the context of inference on the a collection of means, selected on the basis of data from normal distributions, Cox (1975) showed via a theoretical analysis that the efficiency of sample splitting relative to an exact conditional approach is surprisingly high. In the present context, there is a clearer separation of information in the data relevant to the different phases of analysis, so a greater efficiency gain might be expected, although this has to be counterbalanced by the lack of directionality for the full-sample test. These considerations are probed by simulation in §5 for the normal-theory linear regression model.
2. Assessment of model adequacy
2.1. The co-sufficient manifold
The role of sufficiency in the assessment of model adequacy, implicit in R. A. Fisher’s work, is more easily approached via Barndorff-Nielsen and Cox (1994, p. 29). When the ideas can be operationalised, there are no difficulties associated with a double use of the data for model assessment and parametric inference. For notational simplicity throughout the present §2, we drop the subscript $m$ indexing the particular model under consideration. The ideas are at this point detached from the specifics of confidence sets of models and follow the geometric exposition of Battey (2024).
Consider an arbitrary model $\mathcal{E}$, provisionally assumed to contain the true unknown outcome distribution, and having an associated parameter vector $\theta$. The sufficient statistic $S=s(Y)$ for $\theta$ is a function of $Y=(Y_{1},\ldots,Y_{n})$ with observed value $s^{o}=s(y^{o})\in\mathbb{R}^{d}$, where $y^{o}\in\mathbb{R}^{n}$ is the observed value of $Y$. For $n$ independent draws from the same distribution in model $\mathcal{E}$, the joint density or mass function at $y=(y_{1},\ldots,y_{n})$ separates as
$f_{Y}(y;\theta)=\prod_{i=1}^{n}f_{Y_{i}}(y_{i};\theta)=g(s(y);\theta)\prod_{i=1}^{n}h(y_{i})$ (2.1)
by the factorisation theorem (e.g. Barndorff-Nielsen and Cox, 1994, ch. 2.3). An important implication of (2.1), provided that the dimension $d$ of the minimal sufficient statistic $S$ is smaller than $n$, is that there is information in $Y$ not used for inference on $\theta$ and therefore available for assessment of the model.
It is convenient to write the information in $Y$ as $(S,Q(S))$ where, in this notional idealised separation, $Q(S)$ is not a function of $S$ in the conventional sense, and typically does not have an analytic form in terms of the observations; rather $Q(s^{o})$ has the conditional distribution of $Y$, or some one-to-one transformation of $Y$, given $S=s^{o}$. The realisation $s^{o}$ fixes the co-sufficient manifold
$\mathcal{Q}(s^{o})=\{y\in\mathbb{R}^{n}:s(y)=s^{o}\}\subset\mathbb{R}^{n},$ (2.2)
leaving $n-d$ degrees of freedom for variation of $y$ consistent with the constraint $s(y)=s^{o}$, i.e., $\mathcal{Q}(s^{o})$ is a manifold of dimension $n-d$ embedded in $\mathbb{R}^{n}$. Once $s^{o}$ is observed, the notional random variable $Q(s^{o})$ captures the randomness in $Y$ after conditioning on $S=s^{o}$ and is constrained to $\mathcal{Q}(s^{o})$. The implicit general definition of $Q(s^{o})$ is
$\mathop{\mathrm{pr}}(Q(s^{o})\in\mathcal{A})=\mathop{\mathrm{pr}}(Y\in\mathcal{A}|S=s^{o}),$ (2.3)
where $\mathcal{A}\subseteq\mathcal{Q}(s^{o})$ by definition.
Barndorff-Nielsen and Cox (1994, p. 29) encourage an understanding of the separation (2.1) for the assessment of model adequacy by hypothetically treating the data as arising in two stages: in the first stage we learn that $S=s^{o}$ which is used for inference on $\theta$. In the second stage we learn that $Y=y^{o}$, which is effectively one draw from the conditional distribution of $Y$ given $S=s^{o}$. If $y^{o}$ is extreme when calibrated against this conditional distribution, that casts doubt on the adequacy of the model. This is the usual analogue of proof by contradiction underpinning the core Fisherian hypothesis tests. In the context of confidence sets of models, the strategy based on falsification of hypotheses does not involve an alternative hypothesis and thereby evades the post-selection inference problem induced through use of the likelihood ratio statistic with a data-dependent encompassing model.
For some models, notably the linear regression model and canonical generalised linear models, the co-sufficient manifold can be specified explicitly. When this is not the case, it is natural to seek local approximations to the manifold of interest, and since $s(y^{o})=s^{o}$ by construction, it is reasonable to consider the tangent plane to $\mathcal{Q}(s^{o})$ locally at $y^{o}$. Some discussion of this is provided in §6.5. We focus in the present paper on situations for which the co-sufficient manifold is readily deduced.
2.2. The ancillary separation
A second approach to the assessment of model adequacy is available when the dimension $d$ of $S$ is larger than $d_{\theta}$, the dimension of $\theta$. Write $S\cong(S_{1},A)$, where $S_{1}$ estimates $\theta$ and $A$ is an ancillary statistic. Ideally $A$ is ancillary in the strongest sense, having a distribution that does not depend on $\theta$, which implies the factorisation
$f_{S}(s;\theta)=f_{S\mid A}(s\mid a;\theta)f_{A}(a)dtda.$ (2.4)
More generally, definitions of ancillarity in the presence of nuisance parameters are collated on page 38 of Barndorff-Nielsen and Cox (1994). Since the information relevant for inference for $\theta$ is contained in the conditional distribution of $S$ given $A=a^{o}$, the marginal distribution of $A$ is available for model checking.
For the maximal amount of information to be extracted for model assessment, the random variables $A$ and $Q(s^{o})$ should in principle both be sought. Whether this leads to more powerful inference in practice than an approach based on sample splitting, or based on one or other of $A$ and $Q(s^{o})$ individually, is unclear from this abstract discussion because of the possibility of using directed tests when the sample is partitioned, and because the detailed implementation may vary according to the structure of the regression model under consideration. Section 3 gives analytic calculations and methodology for the Gaussian linear regression model.
3. Distributional calculations for Gaussian linear regression
3.1. Distribution on the co-sufficient manifold
Consider a linear regression model in which, under the postulated model, $Y\in\mathbb{R}^{n}$ is normal with mean $\mu=X\theta\in\mathcal{X}$ and covariance matrix $\Sigma=\sigma^{2}I_{n}$. The space $\mathcal{X}$ is the $d_{\theta}$-dimensional linear space spanned by the $d_{\theta}$ columns of $X$, with $(n-d_{\theta})$-dimensional orthogonal complement
$\mathcal{X}^{\perp}=\{v\in\mathbb{R}^{n}:v^{{\rm T}}x=0,\,x\in\mathcal{X}\}\subset\mathbb{R}^{n},$
so that $\mathcal{X}\oplus\mathcal{X}^{\perp}=\mathbb{R}^{n}$.
A minimal sufficient statistic when $\sigma^{2}$ is unknown is either $S=(X^{{\rm T}}Y,\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon})$ if the covariate matrix $X$ is treated as fixed, or $S=(\hat{\theta},\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon},X^{{{\rm T}}}X)$ if $X$ is treated as random, where $\hat{\theta}$ is the ordinary least squares estimator of $\theta$ (also the maximum likelihood estimator) and $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$ is the residual sum of squares. Here we treat $X$ as fixed so that $d=d_{\theta}+1$. The information in the data that is not contained in the sufficient statistic can be extracted through orthogonal projection of $Y$ onto $\mathcal{X}^{\perp}$, followed by an additional conditioning on the realised value of $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$. The first conditioning implicit in the projection restricts to the $(n-d_{\theta})$-dimensional linear subspace $\mathcal{X}^{\perp}$, while the second conditioning restricts the lengths of permissible vectors, so that the co-sufficient manifold is a hypersphere embedded in $\mathcal{X}^{\perp}$, which can also be viewed as an $(n-d)$-dimensional sphere embedded in $\mathbb{R}^{n}$. Depending on context, we will refer to the co-sufficient manifold interchangeably as a hypersphere in $\mathcal{X}^{\perp}$ and an $(n-d)$-dimensional sphere in $\mathbb{R}^{n}$, the term hypersphere meaning that there is one dimension missing from the ambient space. Having restricted to $\mathcal{X}^{\perp}$, we can for most purposes think of $\mathcal{Q}(s^{o})$ as being a unit hypersphere in $\mathbb{R}^{n-d_{\theta}}$, except that the coordinates within the larger space $\mathbb{R}^{n}$ are specified by the orientation of $\mathcal{X}^{\perp}$. The distribution on the sphere is uniform under correct specification of the model, as shown by the following calculation, analogous to ones used in the invariant testing literature (e.g. Kariya, 1980; King, 1980).
The subspaces $\mathcal{X}$ and $\mathcal{X}^{\perp}$ are groups under addition, which implies that for any matrix $U$ with column span $\text{Im}(U)=\mathcal{X}^{\perp}$, the statistic
$q(y,U)=U^{{\rm T}}y/\|U^{{\rm T}}y\|=U^{{\rm T}}y/(y^{{\rm T}}UU^{{\rm T}}y)^{1/2}.$ (3.1)
is invariant under the group with action $g(a,b):y\mapsto ay+b$, where $a>0$, $b\in\mathcal{X}$. The lowest-dimensional specification of $q(y,U)$ containing the same information is obtained by taking the columns of $U$ to be a linear basis for $\mathcal{X}^{\perp}$. The natural choice of $U$ is the $n\times(n-d_{\theta})$ matrix of orthonormal eigenvectors corresponding to the unit eigenvalues of the standard orthogonal projection $M=I_{n}-X(X^{{\rm T}}X)^{-1}X^{{\rm T}}$, the remaining eigenvalues being zero. It follows that $UU^{{\rm T}}=M$ and $U^{{\rm T}}U=I_{n-d_{\theta}}$.
Consider $W=w(Y)=U^{{\rm T}}Y$, which is normally distributed of mean zero and covariance $U^{{\rm T}}\Sigma U=\sigma^{2}I_{n-d_{\theta}}$. In the transformation $w\mapsto(w/\|w\|,w^{{\rm T}}w)=:(q,r^{2})$, the volume element transforms as
$dw_{1}\cdots dw_{n-d_{\theta}}=\frac{(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}\textstyle{\prod_{j=1}^{n-d_{\theta}-1}}dq_{j}}{2\bigl{(}1-\textstyle{\sum_{j-1}^{n-d_{\theta}-1}}q_{j}^{2}\bigr{)}^{1/2}}=\textstyle{\frac{1}{2}}(r^{2})^{\tfrac{n-d_{\theta}}{2}-1}dr^{2}d\text{Vol}^{n-d}(q),$ (3.2)
where
$d\text{Vol}^{n-d}(q)=d\text{Vol}^{n-d_{\theta}-1}(q)=\frac{\textstyle{\prod_{j=1}^{n-d_{\theta}-1}}dq_{j}}{\bigl{(}1-\textstyle{\sum_{j-1}^{n-d_{\theta}-1}}q_{j}^{2}\bigr{)}^{1/2}}$
is the natural volume element on the surface of the hypersphere in $\mathbb{R}^{n-d_{\theta}}$. The joint density function is therefore
$\frac{(r^{2})^{(n-d_{\theta})/2-1}\exp\Bigl{(}-\frac{r^{2}}{2\sigma^{2}}q^{{\rm T}}q\Bigr{)}}{2(2\pi\sigma^{2})^{(n-d_{\theta})/2}}dr^{2}d\text{Vol}^{n-d}(q),$
from which the marginal density function of $Q$ is obtained by integration using a change of variables from $r^{2}$ to $r^{2}q^{{\rm T}}q/2\sigma^{2}$ as
$\frac{\Gamma\bigl{(}\frac{n-d_{\theta}}{2}\bigr{)}(q^{{\rm T}}q)^{-(n-d_{\theta})/2}}{2\pi^{(n-d_{\theta})/2}}d\text{Vol}^{n-d}(q),$ (3.3)
showing that $Q$ is uniformly distributed on the unit hypersphere in $\mathcal{X}^{\perp}$. Since, under the assumed mean model $\mu\in\mathcal{X}$, $R^{2}=W^{{\rm T}}W$ is distributed as $\sigma^{2}$ times a standard $\chi^{2}$ random variable with $n-d_{\theta}$ degrees of freedom, independently of $Q$, the conditional density function of $Y$ given $R^{2}=r^{2}$ is equal to the marginal density function of $Q$, which is explicitly defined in this case.
This calculation can be generalised to allow for variance components in the form $\Sigma=\sigma^{2}(I_{n}+V(\alpha))$ where $V(\alpha)$ is a known matrix function of a vector parameter $\alpha=(\alpha_{1},\ldots,\alpha_{s})^{{\rm T}}$ with $V(0)=0$. In this case, the appropriate analogue of (3.3) is given by equation (10) of Battey and McCullagh (2024). Alternatively, a transformation to uniformity on the unit hypersphere in $\mathcal{X}^{\perp}$ is $q\mapsto A_{\alpha}^{-1/2}q/(q^{{\rm T}}A_{\alpha}^{-1}q)^{1/2}=:Q^{\prime}$, where $\sigma^{2}A_{\alpha}=U^{{\rm T}}\Sigma U$. There are many tests of uniformity on the sphere, whose properties are well established in the literature, by contrast, tests for compatibility with equation (10) of Battey and McCullagh (2024), if available, are not mainstream.
There are some fundamental and operational difficulties in applying the model assessment strategies of §2 directly. These are discussed in §4 where we provide a more elaborate blueprint for implementation of this broad Fisherian program.
3.2. Distribution of the ancillary component
Notice that $R^{2}$ is the residual sum of squares $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$, which is part of the minimal sufficient statistic $S$ and has a distribution independent of $\theta$. It is ancillary for $\theta$ according to the more general of the definitions on page 38 of Barndorff-Nielsen and Cox (1994) and is therefore also available for the assessment of model adequacy via a version of (2.4) allowing for nuisance parameters.
Suppose initially that $\sigma^{2}$ and the mean subspace $\mathcal{X}$ are known, then $R^{2}$ carries all the information relevant for distinguishing, say, normal errors from Student-$t$ errors. This follows from the properties of elliptically symmetric distributions for which the generic form of the density function is
$f_{Y}(y;\mu,\Sigma,g)\propto|\Sigma|^{-1/2}g\bigl{(}(y-\mu)^{{\rm T}}\Sigma^{-1}(y-\mu)\bigr{)},$
the generator $g:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ identifying the member of the elliptically symmetric class. To provide insight, and a basis for some arguments later in the paper, we briefly consider this setting.
Let $U$ be as defined in §3.1. Then $W=U^{{\rm T}}Y$ is a member of the elliptically symmetric class of distributions in $n-d_{\theta}$ dimensions, with zero mean and covariance $\Sigma_{W}=\sigma^{2}I_{n-d_{\theta}}=U^{{\rm T}}\Sigma U$ (e.g. Muirhead, 1982), and therefore has a stochastic representation (Cambanis et al., 1981, Theorem 1) $W\stackrel{{\scriptstyle d}}{{=}}Z\Sigma_{W}^{1/2}Q^{\prime}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution and $Q^{\prime}$ is uniformly distributed on the surface of the unit hypersphere in $\mathbb{R}^{n-d_{\theta}}$. The distribution of $Z$ with density function $f_{Z}$ determines the member of the elliptic class and is related to the generator $g$ for the standardised variable $\Sigma^{-1/2}_{W}W$ as (Fang et al., 1990)
$f_{z}(z)=\frac{2\pi^{(n-d_{\theta})/2}}{\Gamma((n-d_{\theta})/2)}z^{n-d_{\theta}-1}g(z^{2}).$
In the Gaussian case of §3.1, the random variable $R$ satisfies $R\stackrel{{\scriptstyle d}}{{=}}\sigma Z$ where $Z$ is distributed as the square root of a $\chi^{2}$ random variable with $n-d_{\theta}$ degrees of freedom under the null model $\mu\in\mathcal{X}$. As an example, for the Student-$t$ family of errors with $\nu$ degrees of freedom, the distribution of $Z$ is (Kano, 1994) $Z\stackrel{{\scriptstyle d}}{{=}}\surd(\chi^{2}_{n-d_{\theta}}/V)$, where $V\stackrel{{\scriptstyle d}}{{=}}\chi^{2}_{\nu}/\nu$ is independent of the numerator.
The random variable $Z^{2}$ thus carries information on the distribution of the errors when the mean model is known. Conversely, $Z^{2}=R^{2}/\sigma^{2}$ contains valuable information for assessment of the mean model under the assumption that the errors are Gaussian. The practical implementation raises further aspects, to which we return in §4.5.
4. An operational blueprint for Gaussian linear regression
4.1. Some initial observations
From the data, we obtain one realisation $q^{o}$ of the random variable $Q$ from §3.1, which is uniformly distributed on the surface of the unit hypersphere $\mathcal{Q}(s^{o})$ in $\mathcal{X}^{\perp}$. A test based on $q^{o}$ alone has no power in an assessment of model adequacy, as all unit vectors on $\mathcal{Q}(s^{o})$ are compatible with uniformity of distribution. Although power in certain directions could be restored by consideration of alternative models for comparison, this leads to a test statistic essentially equivalent to the $F$-test, i.e. the likelihood ratio test for Gaussian models, and therefore back to the initial difficulties discussed in §1. Note, however, that even with two replicate observations of the random variable $Q$, there would be a degree of power against uniformity of distribution on $\mathcal{Q}(s^{o})$, this tending to increase with the number of replicate observations.
This observation points to an initial suggestion for inference based on co-sufficient information partitioning. Information partitioning, while bearing some superficial similarity to sample splitting, differs in that no observational units are discarded in any of the inferential steps. Since all the data are used for both aspects of inference, reduction and model assessment, no information is relinquished in the reduction to $\hat{\mathcal{E}}$, and virtually no co-sufficient information is lost in the assessment of model adequacy. As it turns out, the simplest partitioning approach, while conceptually revealing, is ineffectual for reasons given in §4.2. We propose in §4.3 a more elaborate reformulation, mirroring to some extent the approach of Rasines and Young (2023) in a different context.
4.2. Preliminary insight: power and naïve data-partitioning
From an encompassing model $\hat{\mathcal{E}}$ computed using all available data, we consider all submodels $\mathcal{E}_{m}$ of reasonable size, seeking to assess each one for compatibility with the data. From $k$ disjoint subsamples of size $n_{k}=n/k$, we consider initially $Q^{(1)},\ldots,Q^{(k)}$, constructed as above but based on disjoint subsamples $(Y^{(1)},X^{(1)}),\ldots,(Y^{(k)},X^{(k)})$.
The resulting co-sufficient manifolds for each subsample, and the corresponding null distributions for $Q^{(1)},\ldots,Q^{(k)}$, while comparable when viewed as a hypersphere embedded in $\mathbb{R}^{n_{k}-d_{\theta}}$, are not comparable when viewed within the larger ambient space $\mathbb{R}^{n_{k}}$, as the column spaces of $X^{(1)},\ldots,X^{(k)}$ define $k$ different partitions $\mathcal{X}^{(i)}\oplus\mathcal{X}^{(i)\perp}=\mathbb{R}^{n_{k}}$ and each co-sufficient manifold is embedded within a different subspace $\mathcal{X}^{(i)\perp}$.
To expose the nature of the difficulties in using these partitioned data to assess compatibility with $\mathcal{E}_{m}$, suppose for some preliminary insight that the null model $\mu=X\theta$ is violated, the true direction of the mean vector being $\mu=X\theta+Z\lambda$. For notational convenience and to avoid repetition, we state the following calculations for a single sample of size $n$, to be reinterpreted later in the context of the subsamples of size $n_{k}$. With $U$, $W$ and $d_{\theta}$ as defined in §3.1, the density function of $W$ at $w$ is
$\frac{\exp\{-(a+w^{{\rm T}}w)/2\sigma^{2}\}}{(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{w^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\Bigr{)}dw,$ (4.1)
where $a=\lambda^{{\rm T}}Z^{{\rm T}}UU^{{\rm T}}Z\lambda$. The transformation $w\mapsto(w/\|w\|,w^{{\rm T}}w)=(q,r^{2})$ gives the joint density function of $Q$ and $R^{2}$ as
$\frac{(r^{2})^{(n-d_{\theta})/2-1}}{2(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{a+r^{2}}{2\sigma^{2}}\Bigr{)}\exp\Bigl{(}-\frac{(r^{2})^{1/2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\Bigr{)}dr^{2}d\text{Vol}^{n-d}(q).$
In the absence of a closed form for the marginal distribution of $Q$, Appendix §A.1 establishes the approximation
$f_{Q}(q)d\text{Vol}^{n-d}(q)\simeq\frac{\Gamma\bigl{(}\frac{n-d_{\theta}}{2}\bigr{)}}{2\pi^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{a}{2\sigma^{2}}\Bigr{)}\exp\biggl{\{}\frac{(n-d_{\theta})q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma\sqrt{2}}\biggr{\}}d\text{Vol}^{n-d}(q),$ (4.2)
which we write as
$f_{Q}(q)d\text{Vol}^{n-d}(q)\simeq\frac{\Gamma\bigl{(}\frac{n-d_{\theta}}{2}\bigr{)}}{2\pi^{(n-d_{\theta})/2}}\cdot h(q)d\text{Vol}^{n-d}(q).$
Approximation (4.2), although obtained under the notional limiting operation $(n-d_{\theta})\rightarrow\infty$, is essentially exact, the only error coming from Stirling’s approximation.
The first multiplicative factor on the right hand side of (4.2) is the density function of $Q$ under the null hypothesis. Thus, power is low when $h(q)$ is close to 1 for all $q$. This occurs when $Z\lambda$ is close to zero or when $Z$ lies close to the column span of $X$. The conclusion agrees qualitatively with the power analysis for the likelihood-ratio test in the absence of any dependencies from reduction, as derived in §4.2 of Lewis and Battey (2025) using a limiting argument in Le Cam-type contiguous alternatives.
Consider now a test of uniformity on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ based on the partitioned data, discarding the orientation of each $\mathcal{X}^{(i)\perp}$ within $\mathbb{R}^{n_{k}}$; $k$ now plays the role of sample size for the null uniform samples. Equation (4.2) suggests that the non-null distribution of $Q^{(i)}$ is well approximated by a von Mises-Fisher distribution with mean direction $m^{(i)}=U^{(i)T}Z^{(i)}\lambda$ and dispersion parameter $\kappa^{(i)}=\sigma^{-1}2^{-1/2}(n_{k}-d_{\theta})\|U^{(i)T}Z^{(i)}\lambda\|$. Since the mean vectors $m^{(i)}$ depend on how the data are partitioned, and because these themselves are in general numerically indistinguishable from uniformly distributed random vectors on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ when standardised to unit length, the distribution of the test statistic when the assumed model is violated is hard to distinguish from that under the model postulated under the null hypothesis. Put differently, by ignoring the orientation of the embedding manifold $\mathcal{X}^{(i)\perp}$, the analysis is made too unconditional.
We propose, instead, the following procedure, which implicitly conditions on the observed column space of $X$, ensuring that pseudo-replicates of $Q$ are constrained to the same co-sufficient manifold as $Q$ when viewed as embedded within $\mathbb{R}^{n-d_{\theta}}$.
4.3. Co-sufficient information partitioning through randomisation
In the vein of Rasines and Young (2023), we first generate independent pseudo-replicates of $Y$ with common distribution $N(\mu,k\sigma^{2}I_{n})$. The idealised but infeasible randomisation procedure is as follows. Let $L$ be an $n\times(k-1)$ matrix, independent of $Y$, with independent $N(0,1)$ entries, and define the $n\times k$-dimensional matrix of pseudo-replicates as
$[\tilde{Y}^{(1)}\cdots\tilde{Y}^{(k)}]=[Y\;L]\Gamma,$ (4.3)
where $\Gamma$ is a $k\times k$ matrix of coefficients tuned to obtain the desired distribution as
$\Gamma=\begin{bmatrix}1&1&1&\cdots&1\\ a_{1}&-b_{1}&-b_{1}&\cdots&-b_{1}\\ 0&a_{2}&-b_{2}&\cdots&-b_{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&a_{k-1}&-b_{k-1}\end{bmatrix}.$ (4.4)
Here $\{a_{i},b_{i}\colon i=1,\ldots,k-1\}$ are deterministic coefficients that can be computed recursively as: $a_{i}=\sigma\tilde{a}_{i}$ and $b_{i}=\sigma\tilde{b}_{i}$, where
$\begin{split}\tilde{a}_{1}=&\;\sqrt{k-1};\hskip 91.04872pt\tilde{b}_{1}=\frac{1}{\tilde{a}_{1}};\\ \tilde{a}_{i+1}=&\;\Bigl{(}k-1-\sum_{j=1}^{i}\tilde{b}_{j}^{2}\Bigr{)}^{1/2};\hskip 27.88358pt\tilde{b}_{i+1}=\frac{1+\sum_{j=1}^{i}\tilde{b}_{j}^{2}}{\tilde{a}_{i+1}}.\end{split}$ (4.5)
Proposition 4.1.
If $Y\sim N(\mu,\sigma^{2}I_{n})$, then $\tilde{Y}^{(1)},\ldots,\tilde{Y}^{(k)}\sim N(\mu,k\sigma^{2}I_{n})$ independently.
A proof is provided in Appendix A.2. The above construction can be generalised to allow for any known positive-definite covariance matrix $\Sigma$ in the distribution of $Y$ by adjusting $\Gamma$ appropriately.
This decomposition of the original response vector yields independent replicates of $Q$, constructed as $\tilde{Q}^{(i)}=U^{{\rm T}}\tilde{Y}^{(i)}/\|U^{{\rm T}}\tilde{Y}^{(i)}\|$ for $i=1,\ldots,k$, where $U$ is that of §3.1. Under correct submodel specification, $\mu=X\theta$, these $k$ replicates are independent and uniformly distributed on the unit hypersphere in $\mathcal{X}^{\perp}$.
For an actionable version of this approach, a reliable estimate of $\sigma^{2}$ is needed. The classical estimator, constructed from the residual sum of squares of the linear model including all covariates in the encompassing set $\hat{\mathcal{E}}$, is consistent provided that $n-|\hat{\mathcal{E}}|\to\infty$. Since $\hat{\mathcal{E}}$ tends to overfit to the data in small samples, it typically underestimates $\sigma^{2}$ unless the sample size is large. Small-sample bias is reduced through the use of the refitted cross-fitting procedure of Fan et al. (2012), who provided statistical guarantees, or the minor adaptation of this estimator proposed below, whose distribution theory is simple and convenient for the Gaussian linear model.
The estimator described by Fan et al (2012) partitions the outcome and covariate data into two disjoint subsets, $(Y^{(1)},X^{(1)})$ and $(Y^{(2)},X^{(2)})$, each of size $n/2$. Fan et al. (2012) proposed to fit a conservative variable selection procedure within each subset, such that the true set $\mathcal{E}^{*}$ is contained in both selected sets ^$\hat{\mathcal{E}}^{(1)}$ and ^$\hat{\mathcal{E}}^{(2)}$ with high probability. Let $\mathcal{X}^{(1)}_{2}$ be the space spanned by the columns of $X^{(1)}$ indexed by ^$\hat{\mathcal{E}}^{(2)}$ and let $\mathcal{X}^{(2)}_{1}$ be the space spanned by the columns of $X^{(2)}$ indexed by ^$\hat{\mathcal{E}}^{(1)}$. The projection matrices onto these two subspaces are written $P^{(1)}_{2}$ and $P^{(2)}_{1}$. The refitted cross-validation estimator $\hat{\sigma}^{2}_{\text{rcv}}$ of $\sigma^{2}$ is then given by the average of the two independent estimators
$\hat{\sigma}_{1}^{2}=\frac{\|P^{(1)}_{2}Y^{(1)}\|_{2}^{2}}{n/2-|\hat{\mathcal{E}}^{(2)}|},\quad\quad\hat{\sigma}_{2}^{2}=\frac{{\|P^{(2)}_{1}Y^{(2)}\|_{2}^{2}}}{n/2-|\hat{\mathcal{E}}^{(1)}|}.$
Empirically, the best approach computes this estimate, or the one below, using only $[\gamma n]$ observations for some $10.5<\gamma<$ 1, so that $n$ in the previous display is replaced by $[\gamma n]$. This further limits the dependence between $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ and $U^{T}Y$, resulting in a better calibration of the test statistic under the null distribution. In simulations, we found that $\gamma=0.6$ was a good choice across a broad range of scenarios.
In order to prove the distribution theory of the test statistic in Theorem 4.1, it is convenient to use a slightly modified version of the refitted cross validation estimator. Let $df_{1}=n/2-|\hat{\mathcal{E}}^{(2)}|$ and $df_{2}=n/2-|\hat{\mathcal{E}}^{(1)}|$ and define the modified refitted cross-validation estimator as
$\hat{\sigma}^{2}_{\text{mrcv}}:=\frac{df_{1}\hat{\sigma}^{2}_{1}+df_{2}\hat{\sigma}^{2}_{2}}{df_{1}+df_{2}}.$ (4.6)
Proposition 4.2.
Let $\nu=df_{1}+df_{2}$. Conditional on |^$|\hat{\mathcal{E}}^{(1)}|$ and |^$|\hat{\mathcal{E}}^{(2)}|$, and on the containments ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(1)}$ and ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(2)}$,
$\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}\sim\sigma^{2}\frac{\chi^{2}_{\nu}}{\nu}$
and
$\mathop{\mathrm{pr}}\bigl{(}\bigl{|}\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}/\sigma^{2}-1\bigr{|}>\delta\bigr{)}\leq 2\exp\bigl{\{}-\textstyle{\frac{1}{2}}\bigl{[}(\nu+\delta+1)+\nu\log(\nu/(\delta+1)\bigr{]}\bigr{\}}.$
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(k)}$ be the $k$ columns of the $n\times k$ matrix obtained by replacing $\Gamma$ in (4.3) by $\hat{\Gamma}$, the realisable version of $\Gamma$ with $\hat{\sigma}$ in place of $\sigma$. Our recommended choice of estimator $\hat{\sigma}$ is $\hat{\sigma}_{\text{mrcv}}$ computed on a portion $[0.6n]$ of the data. With $\hat{Q}^{(i)}=U^{{\rm T}}\hat{Y}^{(i)}/\|U^{{\rm T}}\hat{Y}^{(i)}\|$, Proposition 4.3 quantifies the maximum Euclidean distance between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and the idealised versions $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ that treat $\sigma$ as known. The conclusion is in terms of $k$ and the error rate $a_{n}$ of the estimator $\hat{\sigma}$ of $\sigma$.
It should be kept in mind that, although the notation does not reflect this, the construction of both $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ depends on a postulated model $\mathcal{E}_{m}$, inducing a mean subspace, $\mathcal{X}_{m}$ say.
Proposition 4.3.
Let $\hat{\sigma}$ be an estimator satisfying $|\hat{\sigma}-\sigma|=O_{p}(a_{n})$ and take $k\lesssim a_{n}^{-2}$. Under the null hypothesis $\mathcal{X}=\mathcal{X}_{m}$,
$\max_{i\in[k]}\|\hat{Q}^{(i)}-\tilde{Q}^{(i)}\|_{2}=O_{p}(a_{n}k^{1/2}).$
For the refitted and modified refitted cross-validation estimators of error variance, $a_{n}=n^{-1/2}$, which suggests an upper growth rate of $k$ with $n$ of order $k\lesssim n^{1/2}$.
Proposition 4.3, although reassuring, misses an important point. The relevant metric in which to measure differences between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ depends on the way in which they are used to assess model adequacy, which is via their distribution under the null hypothesis. Since the weak topology is weaker than the norm topology, we can expect better results than those reflected by Proposition 4.3. In particular, an intuitive reaction to the discussion of §3.2 is that the distribution of $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ should be be relatively unaffected by estimation of $\sigma$. The discussion to follow, and the proof of Theorem 4.1 verifies this intuition, and clarifies some more subtle points.
In contrast to the simple data-partitioning scheme considered in §4.2, the replicates of $Q$ obtained through randomisation via (4.3) are approximately independent and identically distributed under mean models other than the postulated one $\mu=X\theta$. As a consequence, this decomposition is more amenable than that of §4.2 to standard testing procedures. While there are many tests of uniformity on a sphere, we focus on the high-dimensional Rayleigh test, as its properties are guaranteed under a high-dimensional regime that matches our requirements, and efficient implementations are available in standard software.
In the present context, the Raleigh test is based on the statistic
$R_{n}:=\frac{\sqrt{2(n-d_{\theta})}}{k}\sum_{1\leq i<j\leq k}\left<\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>\rightarrow_{d}N(0,1),$ (4.7)
or rather, its analogue with $\hat{Y}^{(i)}$ in place of $\tilde{Y}^{(i)}$. This test rejects for large values of $R_{n}$, calibrated against the asymptotic null distribution. This is consistent with a direction-free alternative hypothesis: if $\mu\notin\mathcal{X}$, then $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ are approximately a random sample from projected Gaussian with non-zero mean, so small angles among them are indicative of model misspecification. The distributional convergence was ascertained by Paindaveine and Verdebout (2016) in a more general setting than that needed here. Since, under the null hypothesis that the postulated mean model is correct, $U^{{\rm T}}\tilde{Y}^{(i)}$ is normally distributed of mean zero and variance $k\sigma^{2}I_{n-d_{\theta}}$, exact distribution theory for $R_{n}$ can in principle be derived from the calculations of Fisher (1915). We have
$\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}}=\frac{\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},$ (4.8)
and the standardised quantity in the numerator is standard normally distributed in $n-d_{\theta}$ dimensions. It follows that the distribution of
$\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>=\cos\left(\angle\left(\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{k^{1/2}\sigma},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{k^{1/2}\sigma}\right)\right)=:\bar{R},$
is that of the sample correlation coefficient between two standard normally distributed random variables from $n-d_{\theta}$ observation points. In the above display, $\angle(a,b)$ denotes the angle between vectors $a$ and $b$. The density function of the sample correlation coefficient $\bar{R}$ is (Fisher, 1915)
$f_{\bar{R}}(r)=\frac{\Gamma((n-d_{\theta})/2)}{\sqrt{\pi}\Gamma\{(n-d_{\theta}-1)/2\}}(1-r^{2})^{(n-d_{\theta}-3)/2},\quad-1<r<1.$ (4.9)
That the distribution of $R_{n}$ in (4.7) is asymptotically standard normal follows essentially by the central limit theorem on noting that $\mathbb{E}(\bar{R})=0$ and
[…]
$=$ $\;\frac{\Gamma(\tfrac{n-d_{\theta}}{2})}{2\Gamma(\tfrac{n-d_{\theta}}{2}+1)}\simeq\frac{1}{n-d_{\theta}},\quad(n-d_{\theta}\rightarrow\infty),$
where the equality in the second line uses the gamma recursion $\Gamma(3/2)=\sqrt{\pi}/2$, and the asymptotic form comes from Stirling’s approximation $\Gamma(x+a)/\Gamma(x)\simeq x^{a}$ for $a$ fixed and $x$ large. In fact, for the Gaussian case, the asymptotic distribution is essentially exact for $k$ as small as 2, because in (4.9)
$(1-r^{2})^{(n-d_{\theta}-3)/2}\simeq e^{-(n-d_{\theta})r^{2}/2},\quad(r\rightarrow 0),$ (4.10)
which is the non-normalised density function of a Gaussian random variable with mean zero and variance $(n-d_{\theta})^{-1}$.
Theorem 4.1 shows how estimation of $\Gamma$ using the modified refitted cross-validation estimator (4.6) affects the theoretical guarantee (4.7).
Theorem 4.1.
Conditional on |^$|\hat{\mathcal{E}}^{(1)}|=df_{1}$ and |^$|\hat{\mathcal{E}}^{(2)}|=df_{2}$, and on the containments ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(1)}$ and ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(2)}$,
$\hat{R}_{n}:=$ $\;\frac{\sqrt{2(n-d_{\theta})}}{k}\sum_{1\leq i<j\leq k}\biggl{<}\frac{U^{{\rm T}}\hat{Y}^{(i)}}{\lVert U^{{\rm T}}\hat{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\hat{Y}^{(j)}}{\lVert U^{{\rm T}}\hat{Y}^{(j)}\rVert_{2}}\biggr{>}$
=$\stackrel{{\scriptstyle d}}{{=}}$ $\;R_{n}+O_{p}\Bigl{(}\max\Bigl{\{}k\nu^{-1/2}(n-d_{\theta})^{-1},k\nu^{-1}(n-d_{\theta})^{-1/2}\Bigr{\}}\Bigr{)},\quad(n,\nu\rightarrow\infty,),$
where $\nu=df_{1}+df_{2}\asymp n$. It follows from (4.7) that $\hat{R}_{n}\rightarrow_{d}N(0,1)$ provided that $k\lesssim n^{3/2}$.
Remark 4.1.
Although the estimator of error variance looks to play a minimal role in Theorem 4.1, it is important for the proof strategy that the modified cross validation estimator $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ is used.
The error term is minimised with $k=2$ and in view of (4.10), we expect this choice to provide a close distributional approximation under the null model. There may be reasons to prefer larger $k$, particularly if this provides appreciable improvements in power to detect departures from the null model for only a small loss in coverage. Conditioning on |^$|\hat{\mathcal{E}}^{(1)}|$ and |^$|\hat{\mathcal{E}}^{(2)}|$ in the statement of Theorem 4.1 breaks the dependence between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$. In the simulations of §5, where this aspect is implicitly averaged over in the simulation replicates, some mild dependence is introduced, which may contribute to the deterioration in the distribution theory for increasing $k$.
It may be questioned, in view of the discussion preceding Theorem 4.1, why the randomisation of equation (4.3) is needed at all, as rotational invariance of the Haar measure implies that any vector $v$ on a hypersphere in $\mathbb{R}^{n-d_{\theta}}$ gives rise to an angle $\langle Q,v\rangle$ that follows the distribution in (4.9). This approach, like that of §4.2 ignores the orientation of the embedding manifold $\mathcal{X}^{\perp}$ in $\mathbb{R}^{n}$. Thus, while the null distribution theory is correct in an unconditional sense, the analysis is not relevant for detecting departures from the null model.
4.4. A test based on the ancillary component
As discussed in §3.1 and §3.2, $R^{2}=\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$ has a $\sigma^{2}\chi^{2}_{n-d_{\theta}}$ distribution and is ancillary for $\theta$ under the postulated mean model. To exploit this for model assessment, a consistent estimator of $\sigma^{2}$ is needed, as in §4.3. Proposition 4.4 establishes the error incurred by using the $\chi^{2}_{n-d_{\theta}}$ quantiles of the random variable $V:=R^{2}/\sigma^{2}$ to approximate those of $\hat{V}:=R^{2}/\hat{\sigma}^{2}$, on the basis of arbitrary consistent estimator $\hat{\sigma}^{2}$, which needs to be assumed independent of $R^{2}$ for the theoretical derivations. The two estimators $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ and $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ described in §4.3 clearly do not satisfy this independence assumption, and it does not seem possible to define an estimator with this property from the same sample. Nevertheless, our numerical investigations found that $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ and $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ computed on a proportion $\gamma=0.6$ of the data, have small enough dependence to yield reasonably good inferential performance, although slightly miscalibrated in small samples.
Proposition 4.4.
Let $\hat{\sigma}^{2}$ be a consistent estimator of $\sigma^{2}$, independent of $R^{2}$. Then,
$\sup_{v>0}|\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)|=o(1).$
Proof.
The result is a direct application of Polya’s theorem. An alternative derivation with possibly very conservative rates is given in Appendix B via a more explicit statement about the rate of convergence of $\hat{\sigma}^{2}$. ∎
4.5. On the combination of ancillary and co-sufficient information
In principle, a combination of the ancillary and co-sufficient components achieves the maximal extraction of information relevant for the assessment of model adequacy. In the Gaussian case under consideration here, and under the model postulated for a potential contradiction, a weak form of ancillarity for $\theta$ is used, in which dependence on the nuisance parameter $\sigma^{2}$ is allowed. This dependence, while less critical in conventional uses of ancillarity, is consequential for present purposes, as explained by the following observations.
The random vector $W$ contains both the co-sufficient and the ancillary information. Indeed, when $\sigma^{2}$ is known, there is no ancillary component and $W$ is co-sufficient. Thus, when $\sigma^{2}$ is known, tests of model adequacy would be based on $W$ via $\|W\|^{2}=R^{2}\stackrel{{\scriptstyle d}}{{=}}\sigma^{2}\chi^{2}_{n-d_{\theta}}$. The situation when $\sigma^{2}$ is unknown is more subtle. Although $R^{2}$ and $Q$ are independent, $R^{2}$ is not independent of $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$, the pseudo-replicates of $Q$ constructed with randomisation, which invalidates the use of Fisher’s combination test. Combination tests allowing for dependence are conservative, compromising the calibration accuracy that has been sought throughout the paper, and partially negating the effects of information recovery through combination. The situation is exacerbated once additional dependence is introduced through the estimation of $\sigma^{2}$.
As expected in view of the previous observations, we found in simulations that confidence sets of models based on the ancillary component $R^{2}$ have favourable properties over those based on the co-sufficient pseudo-replicates $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ if $\sigma^{2}$ is known, but that calibration is compromised in small samples once $\sigma^{2}$ is estimated. For sufficiently large sample size, the approach based on the ancillary information with $\sigma^{2}$ estimated is well calibrated.
That the test based on $R^{2}$ is sensitive to estimation of $\sigma^{2}$ is not unexpected: $R^{2}$ estimates $\sigma^{2}$ when the submodel being tested is correct, so it is impossible to completely break the dependence between $R^{2}$ and $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ or $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$, as required in the idealised Proposition 4.4. On the other hand, since $Q$ and $R^{2}$ are independent, $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ and $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ are expected to be asymptotically independent. This can also be seen on noting that $\mathbb{E}(W)$ and $\sigma^{2}$ are orthogonal parameters in the sense of Huzurbazar (1950).
5. Numerical comparisons
5.1. Experimental design
This section explores the numerical performance of the approaches based on the co-sufficient and ancillary separations relative to the simpler approach based on sample splitting, which inevitably discards relevant information. We focus on the Gaussian linear model of §3.1. Conditional on all signal variables surviving the reduction phase, all three approaches have the correct theoretical coverage probability for sufficiently large sample size, but the approach based on the ancillary separation is expected to be more sensitive to estimation of the error variance in small samples, as discussed in §4.5. The extent of this effect is uncovered through simulation in idealised settings. The approach based on sample splitting discards information at both the reduction phase and the model assessment phase, and will therefore discard signal variables with higher probability in the initial phase for an encompassing set $\hat{\mathcal{E}}$ of given size, or conversely, include more variables in $\hat{\mathcal{E}}$ for a fixed probability $\mathop{\mathrm{pr}}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$.
In 500 simulation replicates, $n$ independent rows of $X$ were generated from a normal distribution of dimension $p$ with zero mean and covariance matrix $\Sigma$ having a Toeplitz structure $\Sigma_{ij}=\rho^{|i-j|}$, for $\rho\in(0,1)$. The outcome $Y$ was generated as $Y\sim N(X\theta,\sigma^{2}I_{n})$, where $\sigma^{2}=1$ and $\theta=(t,t,t,0,\ldots,0)^{{\rm T}}$; in other words, the true model is =$\mathcal{E}^{*}=\{1,2,3\}$.
Since the emphasis of this paper is on the model assessment phase, we simplify the reduction phase by choosing $p=400$, so that a single-stage version of Cox reduction can be used rather than the more general two-stage reductions that would be required with a considerably larger number of variables. If $p$ is larger, the reduction phase is more difficult, and the true set of signal variables are less likely in a reduced data set to survive reduction, strengthening the argument for a version that uses all data for both phases of analysis. Detailed analyses of Cox reduction, with extensive numerical comparisons for higher dimensions and different modelling frameworks, are available in Battey and Cox (2018), Hoeltgebaum and Battey (2019) and Lewis and Battey (2025). The latter paper provides insights into how the procedure used at the reduction phase interacts with the model assessment phase. Cox reduction for $p=400$ proceeds by arranging the variable indices at random in a $20\times 20$ grid and traversing the rows and columns, fitting a linear regression to each set of variables. Thus each variable is a covariate in two regressions, alongside different companion variables. Variables that are not significant at a given significance level in either analysis are discarded, the resulting set of variables forming the comprehensive model $\hat{\mathcal{E}}$. In a real application, the significance level in each 20-dimensional regression would ideally be chosen to give an acceptable level of stability for $\hat{\mathcal{E}}$ upon re-randomisation of the variable indices in the grid. For simplicity here, we start with a significance level of $0.05$, reducing it in increments of $0.001$ until the encompassing model $\hat{\mathcal{E}}$ contains at most 15 variables.
For comparison, we also use for reduction an undertuned lasso, with parameter chosen such that approximately the same number of variables are retained in $\hat{\mathcal{E}}$ as were retained by Cox reduction in each simulation replicate. Specifically, the lasso was implemented with the smallest penalty that ensured selection of at most 15 predictors.
Prior to the model-assessment phase, we estimated the error variance via $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ as defined in (4.6), as both the co-sufficiency-based and ancillary-based tests require a consistent estimator of $\sigma^{2}$. This estimator was computed using the first $[0.6n]$ observations, in order to reduce the dependence between the estimator and the test statistics in the idealised known-variance setting. Estimation with the full sample led to invalid tests under the null, whereas the 60% version was found to work well over many scenarios. The conservative variable selection procedure employed for screening in the definition of $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ was the undertuned lasso procedure described before. The lasso, having excellent predictive properties, delivers good estimates of variance even when signal variables are discarded in favour of other variables.
For the construction of the confidence set of models $\mathcal{M}$ we compared the following testing procedures, assessing all subsets of $\hat{\mathcal{E}}$ of size 5 or smaller for their compatibility with the data. Those tests not rejected at the $\alpha=0.05$ significance level constitute $\mathcal{M}$, a $\alpha$-level confidence set of models.
  • Co-sufficient information: $k$ replicates $\hat{Q}_{1},\ldots,\hat{Q}_{k}$ of the spherical vector $Q$ were computed via randomisation as described in §4.3, and were tested for uniformity on the hypersphere in $n-d_{\theta}$ dimensions using the high-dimensional Rayleigh test.
  • Ancillary component: the observed test statistic $\hat{V}=R^{2}/\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ discussed in §4.4 was calibrated against its approximate null distribution $\chi^{2}_{n-d_{\theta}}$.
  • No adjustment for selection: for comparison, we compare the above constructions to the naïve approach of using the same data for reduction and for model assessment based on a likelihood ratio test of each submodel against the comprehensive model. Since we are in a Gaussian setting, the likelihood-ratio statistic reduces to the $F$ statistic. Under each null model $\mathcal{E}_{m}$ to be assessed, this $F$ statistic is calibrated against an $F$ distribution on |^$|\hat{\mathcal{E}}|-|\mathcal{E}_{m}|$ and $n-|\hat{\mathcal{E}}|$ degrees of freedom, which would be the exact distribution under the null model, had $\hat{\mathcal{E}}$ not been selected in the light of the data.
  • Sample splitting: the same $F$ test is performed as described above, except that the first 60% of observations are used for reduction and the remaining 40% for model assessment. This ratio is in the range suggested by Lewis and Battey (2025).
We did not attempt the Markov Chain Monte Carlo approaches noted in the introduction, as there were 4943 submodels of $\hat{\mathcal{E}}$ to be assessed in each simulation run when restricting to models of maximal size 5; MCMC methods seem poorly suited for the confidence sets of models problem in view of the computational requirements.
5.2. Coverage and expected size of the confidence sets
The quantities of interest are reported in Tables 44 for values of $n$, $t$ and $\rho$ each taken at one of two levels in combinations corresponding to a half-replicate of a factorial design: (low, low, high), (high, low, low), (low, high, low), (high, high, high). We also report in Table 5 part of the complementary half-replicate of the full factorial experiment, with each factor at its low level, as this illustrates more striking differences in coverage probabilities from using the co-sufficient or ancillary tests relative to the split-sample $F$ test. These differences are due to the low probability of survival of the full set of signal variables through the reduction phase.
The reported outcome variables are: the simulated probability that $\mathcal{E}^{*}$ is retained through the initial reduction; the simulated coverage probability of the confidence set $\mathcal{M}$; and the simulation average of the number of models in the confidence set. Numbers in parentheses indicate empirical standard deviations of the simulation averages.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{P}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$ $\mathbb{E}(|\mathcal{M}|)$
Cox + co-sufficient ($k=2$) 0.96 (0.01) 1.00 (0.00) 2402 (67)
Cox + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 1674 (60)
Cox + ancillary 0.86 (0.02) 1.00 (0.00) 1574 (59)
Cox + $F$ test 0.08 (0.01) 1.00 (0.00) 17 (2)
Lasso + co-sufficient ($k=2$) 0.94 (0.01) 0.99 (0.01) 2614 (61)
Lasso + co-sufficient ($k=8$) 0.86 (0.01) 0.99 (0.00) 1904 (57)
Lasso + ancillary 0.85 (0.02) 0.99 (0.00) 1821 (57)
Lasso + $F$ test 0.00 (0.00) 0.99 (0.00) 0 (0)
split Cox + split $F$ test 0.92 (0.01) 0.96 (0.01) 1209 (43)
split Lasso + split $F$ test 0.86 (0.02) 0.91 (0.01) 1441 (46)
Table 1. $n=100$; $t=0.5$; $\rho=0.5$.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{P}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$ $\mathbb{E}(|\mathcal{M}|)$
Cox + co-sufficient ($k=2$) 0.99 (0.01) 1.00 (0.00) 2624 (70)
Cox + co-sufficient ($k=8$) 0.96 (0.01) 1.00 (0.00) 1821 (69)
Cox + ancillary 0.95 (0.02) 1.00 (0.00) 1700 (68)
Cox + $F$ test 0.00 (0.00) 1.00 (0.00) 1 (0)
Lasso + co-sufficient ($k=2$) 0.99 (0.01) 1.00 (0.00) 2640 (67)
Lasso + co-sufficient ($k=8$) 0.95 (0.01) 1.00 (0.00) 1865 (67)
Lasso + ancillary 0.95 (0.02) 1.00 (0.00) 1749 (65)
Lasso + $F$ test 0.00 (0.00) 1.00 (0.00) 0 (0)
split Cox + split $F$ test 0.83 (0.02) 0.88 (0.01) 499 (27)
split Lasso + split $F$ test 0.91 (0.01) 0.97 (0.01) 490 (23)
Table 2. $n=150$; $t=0.5$; $\rho=0.1$.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{P}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$ $\mathbb{E}(|\mathcal{M}|)$
Cox + co-sufficient ($k=2$) 0.95 (0.01) 1.00 (0.00) 745 (51)
Cox + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 368 (34)
Cox + ancillary 0.87 (0.01) 1.00 (0.00) 334 (31)
Cox + $F$ test 0.36 (0.02) 1.00 (0.00) 27 (1)
Lasso + co-sufficient ($k=2$) 0.95 (0.01) 1.00 (0.00) 779 (47)
Lasso + co-sufficient ($k=8$) 0.88 (0.01) 1.00 (0.00) 393 (31)
Lasso + ancillary 0.88 (0.01) 1.00 (0.00) 362 (30)
Lasso + $F$ test 0.00 (0.00) 1.00 (0.00) 0 (0)
split Cox + split $F$ test 0.87 (0.01) 0.91 (0.01) 136 (10)
split Lasso + split $F$ test 0.96 (0.01) 0.97 (0.01) 132 (7)
Table 3. $n=100$; $t=1$; $\rho=0.1$.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{P}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$ $\mathbb{E}(|\mathcal{M}|)$
Cox + co-sufficient ($k=2$) 0.97 (0.01) 1.00 (0.00) 270 (13)
Cox + co-sufficient ($k=8$) 0.87 (0.01) 1.00 (0.00) 89 (5)
Cox + ancillary 0.86 (0.02) 1.00 (0.00) 80 (4)
Cox + $F$ test 0.62 (0.02) 1.00 (0.00) 44 (1)
Lasso + co-sufficient ($k=2$) 0.97 (0.01) 1.00 (0.00) 375 (14)
Lasso + co-sufficient ($k=8$) 0.89 (0.01) 1.00 (0.00) 113 (6)
Lasso + ancillary 0.86 (0.02) 1.00 (0.00) 98 (4)
Lasso + $F$ test 0.00 (0.00) 1.00 (0.00) 0 (0)
split Cox + split $F$ test 0.95 (0.01) 1.00 (0.00) 78 (3)
split Lasso + split $F$ test 0.95 (0.01) 1.00 (0.00) 86 (3)
Table 4. $n=150$; $t=1$; $\rho=0.5$.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{P}(\mathcal{E}^{*}\subseteq\hat{\mathcal{E}})$ $\mathbb{E}(|\mathcal{M}|)$
Cox + co-sufficient ($k=2$) 0.91 (0.01) 0.92 (0.01) 3053 (70)
Cox + co-sufficient ($k=8$) 0.87 (0.02) 0.92 (0.01) 2477 (75)
Cox + ancillary 0.86 (0.02) 0.92 (0.01) 2396 (75)
Cox + $F$ test 0.02 (0.01) 0.92 (0.01) 4 (1)
Lasso + co-sufficient ($k=2$) 0.97 (0.01) 0.98 (0.01) 3347 (66)
Lasso + co-sufficient ($k=8$) 0.93 (0.01) 0.98 (0.01) 2750 (75)
Lasso + ancillary 0.92 (0.01) 0.98 (0.01) 2681 (76)
Lasso + $F$ test 0.00 (0.00) 0.98 (0.01) 0 (0)
split Cox + split $F$ test 0.49 (0.02) 0.51 (0.02) 1379 (59)
split Lasso + split $F$ test 0.75 (0.02) 0.79 (0.02) 1532 (62)
Table 5. $n=100$; $t=0.5$; $\rho=0.1$.
The half-replicate allows estimation of the marginal effects of increasing each of $n$, $t$ and $\rho$ from its low to its high level. These are presented in Table 6, where the reported effect on the coverage probability is multiplicative on the odds scale, with values close to one indicative of a null effect, and the reported effect on the expected size is additive on the log scale, with values close to zero suggesting a null effect. Some caution is needed in interpretation of the multiplicative marginal effects, as logistic regression is non-collapsible.
$\mathbb{P}(\mathcal{E}^{*}\subseteq\mathcal{M})$ $\mathbb{E}(|\mathcal{M}|)$
$n$ $t$ $\rho$ $n$ $t$ $\rho$
Cox + co-sufficient ($k=2$) 2.65 0.51 0.64 -0.46 -1.72 -0.55
Cox + co-sufficient ($k=8$) 1.73 0.53 0.53 -0.67 -2.27 -0.75
Cox + ancillary 1.68 0.57 0.55 -0.68 -2.30 -0.75
Cox + $F$ test -1.17 2.12 1.66
Lasso + co-sufficient ($k=2$) 3.28 0.63 0.52 -0.36 -1.58 -0.37
Lasso + co-sufficient ($k=8$) 1.85 0.71 0.60 -0.63 -2.19 -0.61
Lasso + ancillary 1.68 0.65 0.50 -0.67 -2.25 -0.63
Lasso + $F$ test
split Cox + split $F$ test 1.10 1.53 2.59 -0.72 -2.02 0.16
split Lasso + split $F$ test 1.76 2.71 0.69 -0.75 -2.06 0.33
Table 6. Point estimates of marginal effects of $n$, $t$ and $\rho$.
The tables reveal several aspects that were anticipated qualitatively. Use of the $F$ test without sample splitting leads to extremely poor calibration of the confidence set of models, which is restored through sample splitting provided that the true model survives the reduction phase, this being more likely when $n$ and $t$ are large. The role of $\rho$ in the probability of survival is different in the two split-sample versions, high $\rho$ being beneficial for Cox reduction and detrimental for the lasso. Conditional on survival, the co-sufficient and ancillary tests are more conservative in settings where $t$ is low or $\rho$ is high. Additional simulations, not reported, found that the co-sufficient and ancillary tests were exactly calibrated at their nominal levels across all settings when the sample size was sufficiently large or when the true error variance was used instead of $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$, suggesting that the small-sample over-coverage in low-information settings is due to variance estimation. In high-information settings, taking $k$ as large as 20 gives coverage close to nominal, and produces smaller confidence sets than the co-sufficient test with smaller values of $k$, verifying intuition. These high-information settings are, however, cases where the split $F$-test dominates the rest.
In the Gaussian context, because the $F$ test is exact, the coverage probability is only affected by the reduced survival probability in the first stage and not by the small sample size at the second stage, and the efficacy of sample splitting is relatively high, the directionality of the $F$ test compensating to a large extent for the smaller sample size, provided that all signal variables survive reduction. In broad terms, therefore, the main aspect influencing the relative advantage of the co-sufficient and ancillary tests over the split-sample tests is the difficulty of the reduction phase. In non-linear regression models, where a split likelihood ratio test has to be used, the coverage probability of $\mathcal{M}$ is also affected by the smaller sample size in the second stage, due to the $\chi^{2}$ approximation to the likelihood-ratio statistic, although this too could be improved by higher-order correction. We discuss briefly in §6 the extension of co-sufficient model assessment to settings other than the Gaussian linear model.
Between the two values of $k$ reported, intermediate choices of $k$ interpolate between the reported values of Tables 45 in terms of coverage probability and the expected size of the confidence set, with small $k$ achieving higher coverage at the expense of larger confidence sets.
The strategy for reduction is not the focus of this paper. Nevertheless, the tables reflect aspects of the discussion of Lewis and Battey (2025). In particular, Cox reduction leads, on average, to slightly smaller confidence sets than the lasso for the same coverage, as the covariates retained by Cox reduction typically span a larger space than those retained by the lasso. As a result, the retained covariates from Cox reduction fit the data better, making it harder for small sub-models to pass the $F$-test. Cox reduction is also relatively better at retaining signal variables through the reduction phase when the correlation between signal and noise variables is high, and vice versa, as the lasso is more likely to select noise variables that are highly correlated with signal variables.
In all correctly calibrated cases, we see that the size of the confidence set of models is rather large. The maximal size of the confidence set obtained by assessing models up to size 5 from a reduced set with |^$|\hat{\mathcal{E}}|=15$ is 4943. Thus, all approaches have power to reject false models in these low-information settings. Cox and Battey (2017) discuss how to extract compact messages from large confidence sets of models.
6. Extensions to other settings
6.1. Discussion and immediate extension
The ideas of §2.1 and §2.2 apply in principle to any parametric regression model in which, under the model hypothesis, a sufficient statistic or approximate sufficient statistic of smaller dimension than $n$ is available. An exact or approximate analytic approach for non-Gaussian models is likely to lead to more marked improvements over sample splitting relative to the Gaussian case, as the $F$ test for model adequacy is exact, conditional on the true model $\mathcal{E}^{*}$ having survived the first stage, while the log-likelihood test needed for more general models models is based on an asymptotic approximation.
The details of implementation for non-Gaussian models need to be worked out on a case by case basis and may well involve considerable methodological innovation beyond the ideas presented in §4. An exception is the use of the log-normal distribution for modelling positive outcomes, in which case the ideas of §4 apply directly after logarithmic transformation. The log-normal distribution is one of several popular parametric distributions for survival outcomes with covariate-dependence usually modelled in traditional accelerated-life form.
The extension to other models is less direct. In §6.2 and §6.4 we emphasise the structure in canonical exponential-family regression models and in accelerated life models that simplifies the required calculations. The logistic regression case is developed in a little more detail in §6.3, although without rigorous justification.
6.2. Canonical exponential family regression models
An important role is played by canonical exponential family regression models for outcomes $Y=(Y_{1},\ldots,Y_{n})$, assumed conditionally independent given observations on covariates $x$. Assuming the model is correctly specified, the conditional probability density or mass function of $Y=(Y_{1},\ldots,Y_{n})$ at $y$ is
$f(y;x_{1}^{{\rm T}}\theta,\ldots,x_{n}^{{\rm T}}\theta)=\exp\Bigl{[}\phi^{-1}\Bigl{\{}\theta^{T}\sum_{i=1}^{n}x_{i}y_{i}-\sum_{i=1}^{n}K(x_{i}^{T}\theta)\Bigr{\}}\Bigr{]}\prod_{i=1}^{n}h(y_{i},\phi^{-1}),$
where $\phi$ is a dispersion parameter, always equal to 1 in some models. Example cases are $K(\eta)=\eta^{2}/2$ for the normal theory linear model, $K(\eta)=e^{\eta}$ for Poisson regression, and $K(\eta)=\log(1+e^{\eta})$ for binary regression.
Provided that any floating dispersion parameter is known, the sufficient statistic for $\theta$ is $S=\sum_{i=1}^{n}x_{i}Y_{i}=X^{{\rm T}}Y$, where each $x_{i}$ is of the same dimension $d=d_{\theta}$ as $\theta$. There is a reduction by sufficiency from $n$ to $d$ and no ancillary statistic in the sense of §2.2. The insights of §6.5 to be presented show that the co-sufficient manifold $\mathcal{Q}(s^{o})$ is flat. We consider in more detail the example of logistic regression.
6.3. Logistic regression
In contrast to §3, where the conditional distribution of interest reduced to a marginal distribution by the properties of Gaussian distributions, the co-sufficient statistic $Q(s^{o})$ is notional in the sense that it cannot be expressed in terms of observable quantities and is given implicitly by the defining equation (2.3). The conditional distribution of interest is
$\mathop{\mathrm{pr}}(Y=y\mid S=s^{o})=\frac{\mathop{\mathrm{pr}}(Y=y)}{\sum_{y^{\prime}\in\mathcal{Q}(s^{o})}\mathop{\mathrm{pr}}(Y=y^{\prime})}.$ (6.1)
With a different objective in mind, Hanley and Roy (2025) obtained by linear integer programming, all sets of binary outcomes $y\in\{0,1\}^{n}$ compatible with the observed value of the sufficient statistic $s^{o}=X^{{\rm T}}y^{o}$, i.e. a complete enumeration of $\mathcal{Q}(s^{o})$, which is a finite set in this case. Their original motivation for obtaining these solutions was to re-analyse by a superficially more modern approach, an old analysis due to Penrose (1933), aided by Fisher. The conclusion of their pedagogical piece was that all solutions to this linear integer programming problem gave the same inference for the logistic regression parameters as that obtained by Penrose and Fisher, as they must by construction, by the definition of sufficiency. Summation over Hanley’s and Roy’s (2025) full solution set in equation (6.1) provides a means of calculating the conditional distribution of interest. In practice, the probabilities on the right hand side of (6.1) depend on the unknown parameter vector $\theta$, which can be replaced by the maximum likelihood solution $\hat{\theta}$.
6.4. Accelerated life models
Let $T_{0}$ be a positive random variable representing the outcome of individuals with covariates at baseline $z=0$. Here we use $z$ in place of $x$, as $x$ is defined to be one dimension larger and includes a column of ones. The density function transforms under scaling as
$f_{\sigma T_{0}}(t)dt=\sigma^{-1}f_{T_{0}}(t/\sigma)dt,$
implying that $T=T_{0}/\rho(z)$ has a density function of accelerated life form
$f_{T}(t;z^{{\rm T}}\beta)=\rho(z)f_{T_{0}}(\rho(z)t),$
where the effect of the covariates is to multiply the baseline rate $\rho_{0}$ by $\rho(z)$, or equivalently to multiply the baseline scale $\sigma_{0}$ by $\sigma(z)$. With $\rho(z)=\exp(z^{{\rm T}}\beta)$,
$Y:=\log T=\log T_{0}-\log\rho(z)=\log T_{0}-z^{{\rm T}}\beta=\mu_{0}-z^{{\rm T}}\beta+\varepsilon,$ (6.2)
where $\mu_{0}=\mathbb{E}(\log T_{0})$ and $\varepsilon$ is a mean-zero error term, whose distribution is determined by that of the baseline variable $T_{0}$. The broad ideas of §3 apply to the log-transformed outcome model (6.2) with $\theta$ one dimension larger than $\beta$ so as to include the intercept $\mu_{0}$. If the baseline outcome distribution is log-normally distributed, no modification is needed. Here we consider an alternative choice of outcome distribution to highlight the considerations involved.
Suppose for concreteness that $T_{0}$ has a log-logistic distribution of rate $\rho_{0}>0$ and scale $\tau>0$, then $Y_{0}=\log T_{0}$ has a logistic distribution of mean $\log\rho_{0}$ and variance $\tau^{2}$ so that $Y$ satisfies (6.2) with $\varepsilon$ having density function
$f_{\varepsilon}(\epsilon)=\frac{\exp(-\epsilon/\tau)}{\tau(1+\exp(-\epsilon/\tau))},$ (6.3)
With a slight ambiguity in notation, resolved by context, $\varepsilon$ denotes both the random variable with density function (6.3) and the vector of independent replicates following the same distribution. These independent replicates are written $\varepsilon_{1},\ldots,\varepsilon_{n}$. Let $X$ be the $n\times d_{\theta}$-dimensional matrix of covariate information. Following §3.1, take $U=(u_{ij})\in\mathcal{X}^{\perp}$ to be the $n\times(n-d_{\theta})$ matrix of orthonormal eigenvectors corresponding to the unit eigenvalues of $M=I_{n}-X(X^{{\rm T}}X)^{-1}X^{{\rm T}}$ and let $W=U^{{\rm T}}Y$. Under the null model, $\mu=X\theta\in\mathcal{X}$, and the distribution of $W$ does not depend on $\theta$.
In view of the discussion of §4.5, we only consider inference based on $W$, rather than on $Q=U^{{\rm T}}Y/\|U^{{\rm T}}Y\|$. The latter has a distribution on the unit hypersphere in $\mathcal{X}^{\perp}$, however this distribution is not uniform and is more difficult to calculate in the present setting. The characteristic function $\varphi_{W}(z)$ of $W$ can be written in terms of the characteristic function $\varphi_{\varepsilon}$ of $\varepsilon$ as $\varphi_{W}(z)=\varphi_{\varepsilon}(U^{{\rm T}}z)$ where
$\varphi_{\varepsilon}(t)=\prod_{i=1}^{n}\varphi_{\varepsilon_{i}}(t_{i})=(\pi\tau)^{n}\prod_{i=1}^{n}\frac{t_{i}}{\sinh(\pi\tau t_{i})}.$ (6.4)
Inversion of $\varphi_{W}(z)$ gives the density function
[…]
where $\Re(\cdot)$ denotes the real part of a complex number. Voldoire et al. (2024) suggest a procedure for obtaining samples from (6.5). By comparing observed values of $W$ to the tails of this sampling-based distribution, adequacy of the model can be assessed. This requires the logistic standard deviation $\tau$ in (6.4) to be replaced by a consistent estimate, as in §4.3. Ideally, the integral (6.5) might instead be approximated analytically, as sampling-based approaches are prohibitively expensive in terms of computation for reasons mentioned earlier.
6.5. General non-exponential family models
The $n\times d$ matrix of gradient vectors
$V(y^{o})=\left.\frac{\partial s^{{\rm T}}(y)}{\partial y}\right|_{y=y^{o}}$ (6.6)
specify the normal directions to $\mathcal{Q}(s^{o})$ at $y^{o}$. The space spanned by the columns of $V(y^{o})$ is the normal subspace of $\mathcal{Q}(s^{o})$ at $y^{o}$ and the orthogonal complement
$V^{\perp}(y^{o})=\{u\in\mathbb{R}^{n}:u^{{\rm T}}V(y^{o})=0\in\mathbb{R}^{d}\}$
defines the tangent plane at $y^{o}$. For canonical exponential-family regression models with no floating dispersion parameter these matrices do not depend for $y^{o}$, implying that the co-sufficient manifold is flat. This idea underpins the tangent exponential model of Fraser (1988, 1990) and Fraser and Reid (1988, 1995), whose objective was higher order conditional inference on $\theta$. Our problem is in essence the opposite of theirs. The establishment of general analytic approximations is an open problem. In principle, such approximations could replace the importance-sampling based approach suggested for the log-logistic accelerated life model of the previous section, and possibly also the integer-programming based approach for the logistic regression formulation of §6.3.
7. Closing discussion
The literature on post-selection inference has focused on the conceptually different problem of performing inference on a parameter following a data-based selection of effects on which formal inference is sought. The problem of post-reduction inference for confidence sets of models raises similar challenges, but is conceptually very different. In particular, the object of inference in the confidence sets of models problem is unambiguously defined, and the two phases of analysis are such that there is a clearer separation of information in the data for the different phases. This contrasts strongly with the more widely studied post-selection inference problems, which address almost the same scientific question in different guises: a selection phase asking whether the effect is zero, and an inference phase asking what values of strongly-suggested effects are compatible with the data.
In the confidence sets of models formulation, the most relevant hypotheses for test, as small submodels of $\hat{\mathcal{E}}$, are stable over hypothetical replication from the same generating process. Provided that $\hat{\mathcal{E}}$ reliably includes $\mathcal{E}^{*}$, the double use of the data is only problematic if the overfitted $\hat{\mathcal{E}}$ is used as a basis for comparison of fit, as in the likelihood ratio test. The present paper showed how to avoid this comparison by operationalising the appropriate co-sufficiency and ancillary separations. If the ancillary statistic for the interest component depends weakly on nuisance parameters as in the weaker definitions of ancillarity on page 38 of Barndorff-Nielsen and Cox, then the small-sample coverage properties of the ancillary-based assessment may be compromised through estimation of these nuisance parameters. Numerical work showed that, although the efficiency of sample splitting is relatively high for the Gaussian model under consideration here due to the exactness of the second-stage test, there are modest advantages of co-sufficient information partitioning in settings where the reduction phase is difficult, for instance when the sample size is small relative to dimension or the signal strength is low.
If the modelling assumptions are accepted, the exact analytic calculations of the present paper provide insight, reduce approximation errors, and considerably reduce computation relative to approaches based on Monte Carlo sampling.
A natural first reaction to the tables in §5 is that the number of models in the confidence set is inconveniently large. This is an honest reflection of the limitations of the data, and emphasises the dangers of reporting a single model when many are compatible with the data at the chosen significance level. Statistics can take us no further, and any choice between models in the confidence set requires either additional data or subject-matter expertise. Cox and Battey (2017) discuss how one might extract compact messages from large confidence sets of models. Specifically, among the large number of models reported in their supplementary material, two variables, call them $v_{1}$ and $v_{2}$, are present in 96% and 94% of cases. In 78% of the models in which $v_{2}$ is not present, another variable, $v_{3}$, is present in its place, and only 1% of models include neither $v_{2}$ nor $v_{3}$.
If prediction is the sole objective, an arbitrary well-fitting model, such as that selected by the lasso, will be equally effective and there is no need for a confidence set of models, although they could be used to provide a confidence distribution for predictions as discussed in Section S3 of the supplement of Lewis and Battey (2025).
Acknowledgements
HB is grateful to James Hanley for sharing a preprint of Hanley and Roy (2025). EPSRC fellowship EP/T01864X/1 is gratefully acknowledged.
SUPPLEMENTARY MATERIAL
Appendix A Proofs and derivations
A.1. Derivation of equation (4.2)
From the joint density function
$\frac{(r^{2})^{(n-d_{\theta})/2-1}}{2(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{a+r^{2}}{2\sigma^{2}}\Bigr{)}\exp\Bigl{(}-\frac{(r^{2})^{1/2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\Bigr{)}dr^{2}d\text{Vol}^{n-d}(q)$
of $Q$ and $R^{2}$, expand the exponential function around zero in $\exp(-(r^{2})^{1/2}q^{{\rm T}}U^{{\rm T}}Z\lambda/\sigma^{2})$. This leads to an infinite-series representation of the marginal density function of $Q$:
[…]
After a change of variables from $r^{2}$ to $r^{2}/2\sigma^{2}$, this is solved as the infinite series of gamma integrals
$f_{Q}(q)d\text{Vol}^{n-d}(q)=\frac{\exp(-a/2\sigma^{2})}{2\pi^{(n-d_{\theta})/2}}\sum_{m=0}^{\infty}\frac{1}{m!}\Bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\Bigr{)}^{m}\Gamma\Bigl{(}\frac{n-d_{\theta}+m}{2}\Bigr{)}d\text{Vol}^{n-d}(q),$ (A.1)
which is shown to be convergent by consideration of the ratios of successive terms:
$M_{m}=\frac{m!\bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\bigr{)}^{m+1}\Gamma\bigl{(}\frac{n-d_{\theta}+(m+1)}{2}\bigr{)}}{(m+1)!\bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\bigr{)}^{m}\Gamma\bigl{(}\frac{n-d_{\theta}+m}{2}\bigr{)}}=\Bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{(m+1)\sigma}\Bigr{)}\frac{\Gamma\bigl{(}\frac{n-d_{\theta}+(m+1)}{2}\bigr{)}}{\Gamma\bigl{(}\frac{n-d_{\theta}+m}{2}\bigr{)}}.$
On using Stirling’s asymptotic approximation to the gamma integral in the form $\Gamma(x+\alpha)/\Gamma(x)\simeq x^{\alpha}$ for large $x$ and $\alpha$ fixed we see that $\lim_{m\rightarrow\infty}M_{m}=0$, so that the series representation for $f_{Q}(q)d\text{Vol}^{n-d}(q)$ is convergent. On multiplying and dividing (A.1) by $\Gamma\{(n-d_{\theta})/2\}$ and approximating the resulting gamma ratios using Stirling’s approximation as above, we obtain the approximation
$f_{Q}(q)d\text{Vol}^{n-d}(q)$ $\simeq$ $\frac{\exp(-a/2\sigma^{2})}{A}\sum_{m=0}^{\infty}\frac{1}{m!}\Bigl{(}\frac{\sqrt{2}q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma}\Bigr{)}^{m}\Bigl{(}\frac{n-d_{\theta}}{2}\Bigr{)}^{m}d\text{Vol}^{n-d}(q)$
$=$ $\frac{\exp(-a/2\sigma^{2})}{A}\exp\Bigl{(}\frac{(n-d_{\theta})q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma\sqrt{2}}\Bigr{)}d\text{Vol}^{n-d}(q),$
where, on recalling that $d=d_{\theta}+1$,
[…]
is the surface area of the unit hypersphere $\mathcal{S}_{n-d_{\theta}-1}$ embedded in $\mathbb{R}^{n-d_{\theta}}$. This is expression (4.2).
A.2. Proof of Proposition 4.1
Proof.
The $i$-th replicate of $Y$ can be written as $\tilde{Y}^{(i)}=Y-\sum_{j=1}^{i-1}b_{j}L_{j}+a_{i}L_{i}$, where $L_{i}$ is the $i$-th column of $L$, $a_{k}=0$, and empty sums of the form $\sum_{k=1}^{0}$ are defined to be 0. We need to verify that all replicates have covariance $k\sigma^{2}I_{n}$ and are independent. We have:
$\mathrm{Cov}(\tilde{Y}^{(i)})=\Bigl{(}\sigma^{2}+\sum_{j=1}^{i-1}b_{j}^{2}+a_{i}^{2}\Bigr{)}I_{n}=\Bigl{(}\sigma^{2}+\sum_{j=1}^{i-1}b_{j}^{2}+\sigma^{2}(k-1)-\sum_{j=1}^{i-1}b_{j}^{2}\Bigr{)}I_{n}=k\sigma^{2}I_{n}.$
For every $1\leq i_{1}<i_{2}\leq k$,
$\mathrm{Cov}(\tilde{Y}^{(i)},\tilde{Y}^{(j)})=\Bigl{(}\sigma^{2}+\sum_{j=1}^{i_{1}-1}b_{j}^{2}-a_{i_{1}}b_{i_{1}}\Bigr{)}I_{n}=\Bigl{(}\sigma^{2}+\sum_{j=1}^{i_{1}-1}b_{j}^{2}-\sigma^{2}-\sum_{j=1}^{i_{1}-1}b_{j}^{2}\Bigr{)}I_{n}=0_{n\times n},$
where $0_{n\times n}$ is a matrix of 0’s and we are taking $b_{0}=0$ if $i_{1}=1$. ∎
A.3. Proof of Proposition 4.2
Proof.
The distributional statement is standard. Let $T:=\hat{\sigma}^{2}_{\text{mrcv}}/\sigma^{2}$ and $\nu=df_{1}+df_{2}$. The moment-generating function of a $\chi^{2}_{\nu}$ random variable is
$M(q)=(1-2q)^{-\nu/2},\quad q<\textstyle{\frac{1}{2}}.$
Thus, by Chernoff’s inequality, for any $\delta>0$ and $q<\textstyle{\frac{1}{2}}$,
$\mathop{\mathrm{pr}}(T-1>\delta)\leq e^{-q(\delta+1)}(1-2q)^{-\nu/2}.$
On substituting in the minimiser $q=(\delta+1-\nu)/2(\delta+1)$, this bound is
$\mathop{\mathrm{pr}}(T-1>\delta)\leq\exp\Bigl{(}-\frac{\delta+1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}^{-\nu/2}.$
Similarly, for any $\delta\in[0,1)$,
$\mathop{\mathrm{pr}}(T-1<-\delta)\leq\;e^{q(\delta-1)}(1-2q)^{-\nu/2}\leq\;\exp\Bigl{(}\frac{\delta-1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{1-\delta}\Bigr{)}^{-\nu/2}.$
Let
$u(\nu,\delta):=\exp\Bigl{(}-\frac{\delta+1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}^{-\nu/2}.$
Since
$\exp\Bigl{(}\frac{\delta-1+\nu}{2}\Bigr{)}\Bigl{(}\frac{\nu}{1-\delta}\Bigr{)}^{-\nu/2}\leq u(\nu,-\delta),$
we have
$\mathop{\mathrm{pr}}(|T-1|>\delta)\leq 2\exp\Bigl{[}-\frac{1}{2}\Bigl{\{}(\delta+1+\nu)+\nu\log\Bigl{(}\frac{\nu}{\delta+1}\Bigr{)}\Bigr{\}}\Bigr{]}.$
A.4. Proof of Proposition 4.3
Proof.
Since
$\hat{\Gamma}-\Gamma=(\hat{\sigma}-\sigma)\begin{bmatrix}1&0&0&\cdots&0\\ \tilde{a}_{1}&-\tilde{b}_{1}&-\tilde{b}_{1}&\cdots&-\tilde{b}_{1}\\ 0&\tilde{a}_{2}&-\tilde{b}_{2}&\cdots&-\tilde{b}_{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\tilde{a}_{k-1}&-\tilde{b}_{k-1}\end{bmatrix},$
where $\tilde{a}_{i}\asymp k^{1/2}$ and $\tilde{b}_{i}\asymp k^{-1/2}$ are defined in equation (4.5), the components $\hat{Y}^{(1)},\ldots,\hat{Y}^{(k)}$ can be written in terms of $\tilde{Y}^{(1)},\ldots,\tilde{Y}^{(k)}$ as
$\hat{Y}^{(i)}=\tilde{Y}^{(i)}+[Y\;L](\hat{\Gamma}_{i}-\Gamma_{i})=\tilde{Y}^{(i)}+\delta 1_{n}$ (A.2)
for all $i$, where $\delta=O(a_{n}\sqrt{k})$.
Since $U$ has singular values 0 and 1,
$\|\hat{Q}^{(i)}-\tilde{Q}^{(i)}\|_{2}\leq$ $\;\biggl{\|}\frac{\hat{Y}^{(i)}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}-\frac{\tilde{Y}^{(i)}}{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}\biggr{\|}_{2}$
$=$ $\;\biggl{\|}\frac{\delta 1_{n}}{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}-\frac{\hat{Y}^{(i)}}{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}\biggl{(}\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}-1\biggr{)}\biggr{\|}_{2}=O_{p}(a_{n}\sqrt{k}),$
where the last equality follows by Lemma A.1 and because $\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}\asymp(n-d_{\theta})^{1/2}\sqrt{k}$ by orthonormality of the columns of $U$, which implies that the entries are of order $n^{-1/2}$ on average.
Lemma A.1.
Under the conditions of Propisition 4.3,
$\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}=1+O_{p}(a_{n}\sqrt{k}).$
Proof.
By (A.2),
$U^{{\rm T}}\hat{Y}^{(i)}=U^{{\rm T}}\tilde{Y}^{(i)}+U^{{\rm T}}[Y\;L](\hat{\Gamma}_{i}-\Gamma_{i})=U^{{\rm T}}\tilde{Y}^{(i)}+\delta_{u}1_{n-d_{\theta}},$
where $\delta_{u}=O(a_{n}k^{1/2})$ by orthonormality of the columns of $U$, which implies that the entries are of order $n^{-1/2}$ on average. Thus,
$\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}\leq\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\max\{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}-\|\delta_{u}1_{n-d_{\theta}}\|_{2},0\}},$ (A.3)
and it only remains to show that that $\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}-\|\delta_{u}1_{n-d_{\theta}}\|_{2}>0$ on an event of asymptotic probability 1. This follows on noting that $(\sigma k)^{-1}\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}^{2}$ has a $\chi^{2}$ distribution on $n-d_{\theta}$ degrees of freedom. Thus, by Lemma 1 of Laurent and Massart (2000),
$\mathop{\mathrm{pr}}\Bigl{(}(\sigma^{2}k)^{-1}\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}^{2}>(n-d_{\theta})-2(n-d_{\theta})^{1/2}\sqrt{t}\Bigr{)}>1-\exp(-t)$
and consequently
$\mathop{\mathrm{pr}}\biggl{(}\bigcap_{i=1}^{k}\Bigl{\{}\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}^{2}>\sigma k(n-d_{\theta})^{1/2}\Bigl{[}(n-d_{\theta})^{1/2}-2\sqrt{t}\Bigr{]}\Bigr{\}}\biggr{)}>\{1-\exp(-t)\}^{k}$
Thus $\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}\gtrsim a_{n}(k(n-d_{\theta}))^{1/2}$ for all $i$ with probability bounded below by a term of order
$\biggl{\{}1-\exp\biggl{(}-\frac{(n-d_{\theta})a_{n}^{2}}{2k}\biggr{)}\biggr{\}}^{k}\rightarrow 1.$
It follows by a Taylor expansion for small $\|\delta_{u}1_{n-d_{\theta}}\|_{2}=O_{p}(a_{n}k^{1/2})$ in the right-hand side of (A.3) that
$\frac{\|U^{{\rm T}}\tilde{Y}^{(i)}\|_{2}}{\|U^{{\rm T}}\hat{Y}^{(i)}\|_{2}}=1+O_{p}(a_{n}\sqrt{k}).$
A.5. Proof of Theorem 4.1
Proof.
Consider a stochastic expansion of the standardised random variable
$T^{(i)}=U^{{\rm T}}\hat{Y}^{(i)}/(k\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}})^{1/2}.$
Let $\nu=df_{1}+df_{2}$ as given in the definition of the modified refitted cross-validation estimator (4.6). Then by Proposition 4.2 the entries of $T^{(i)}$ have a student-$t$ distribution with $\nu$ degrees of freedom. Let $T$ denote an arbitrary random variable with this same distribution. Its density function at $t$ is
$f_{T}(t)dt=\frac{\Gamma\{(\nu+1)/2\}}{\Gamma(\nu/2)\sqrt{\pi\nu}}(1+t^{2}/\nu)^{-(\nu+1)/2}dt.$
By Stirling’s formula in the form $\Gamma(x+a)/\Gamma(x)\simeq x^{a}$ for $a$ fixed and $x\rightarrow\infty$,
$\frac{\Gamma\{(\nu+1)/2\}}{\Gamma(\nu/2)}\simeq\left(\frac{\nu}{2}\right)^{1/2}\quad(\nu\rightarrow\infty).$
so that the constant term in $f_{T}$ is approximately $(2\pi)^{1/2}$ for large $\nu$. On taking logarithms and expanding for small $t^{2}/\nu$,
$-\frac{(\nu+1)}{2}\log\Bigl{(}1+\frac{t^{2}}{\nu}\Bigr{)}=$ $-\frac{(\nu+1)}{2}\biggl{\{}\sum_{k=1}^{\infty}(-1)^{k+1}\frac{(t^{2}/\nu)^{k}}{k}\biggr{\}}$
$=$ $-\frac{t^{2}}{2}+\nu^{-1}\biggl{(}\frac{t^{4}}{4}-\frac{t^{2}}{2}\biggr{)}+\nu^{-2}\biggl{(}\frac{t^{4}}{4}-\frac{t^{6}}{6}\biggr{)}+O(\nu^{-3}),$
and the density function of $T$ satisfies
$f_{T}(t)=\phi(t)\biggl{\{}1+\nu^{-1}\biggl{(}\frac{t^{4}}{4}-\frac{t^{2}}{2}\biggr{)}+\nu^{-2}\biggl{(}\frac{t^{4}}{4}-\frac{t^{6}}{6}\biggr{)}\biggr{\}}+O(\nu^{-3})$
where $\phi$ is the standard normal density function. It follows that, for $Z$ a standard normal random variable, we can write the stochastic expansion
$T\stackrel{{\scriptstyle d}}{{=}}Z+b_{1}(Z)/\nu^{1/2}+b_{2}(Z)/\nu+O_{p}(\nu^{-3/2}),$ (A.4)
where $b_{1}(z)$ can be obtained from equation (3.73) of Barndorff-Nielsen and Cox (1989) and $b_{2}(z)$ can be obtained from equation (3.72).
Since the above stochastic expansion holds for every entry of $T^{(i)}$ defined above, the entries being asymptotically independent, we can write
$T^{(i)}\stackrel{{\scriptstyle d}}{{=}}Z^{(i)}+B_{1}^{(i)}/\nu^{1/2}+B_{2}^{(i)}/\nu+O_{p}(\nu^{-3/2}),$ (A.5)
where $Z^{(i)}$ is an $(n-d_{\theta})$-dimensional vector of independent standard normal random variables and $B_{1}^{(i)}$ and $B_{2}^{(i)}$ are random variables with entries distributed marginally as $b_{1}(Z)$ and $b_{2}(Z)$ respectively.
We will prove that the entries of the $(n-d_{\theta})$-dimensional vectors $T^{(i)}$ are independent. Suppose for a contradiction that they are not. Then either $Z^{(i)}$ must have dependent entries, or $B_{1}^{(i)}/\nu^{1/2}$ or subsequent terms must have dependent entries. Since $\hat{\sigma}^{2}\rightarrow\sigma^{2}$ as $\nu\rightarrow\infty$, in this limit, $T^{(i)}$ is mean-zero Gaussian with covariance matrix $I_{n-d_{\theta}}$, contradicting dependence in the entries of $Z^{(i)}$. Independence in the entries of $Z^{(i)}$ contradicts independence of the entries of $B_{1}^{(i)}/\nu^{1/2}$, $B_{2}^{(i)}/\nu$, etc. as the stochastic expansion (A.4) shows that these only depend on the corresponding entry of $Z^{(i)}$. That both $\nu$ and the dimension of $T^{(i)}$ are increasing with $n$ does not affect this argument.
Thus write
$\bigl{\langle}T^{(i)},T^{(j)}\bigr{\rangle}\stackrel{{\scriptstyle d}}{{=}}$ $\;\bigl{\langle}Z^{(i)},Z^{(j)}\bigr{\rangle}+\bigl{\langle}Z^{(i)},B_{1}^{(j)}/\nu^{1/2}\bigr{\rangle}+\bigl{\langle}Z^{(j)},B_{1}^{(i)}/\nu^{1/2}\bigr{\rangle}+\cdots$
$=$ $\;\bigl{\langle}Z^{(i)},Z^{(j)}\bigr{\rangle}+O_{p}(\max\{((n-d_{\theta})\nu)^{-1/2},\nu^{-1}\})=O_{p}((n-d_{\theta})^{-1}).$
By the discussion of §3.2, the squared lengths $\|T^{(i)}\|^{2}$ have the distribution of $\chi^{2}_{n-d_{\theta}}/V\stackrel{{\scriptstyle d}}{{=}}(n-d_{\theta})F_{n-d_{\theta},\nu}$ where $V\stackrel{{\scriptstyle d}}{{=}}\chi^{2}_{\nu}\nu$ is independent of the numerator $\chi^{2}_{n-d_{\theta}}$ and $F_{n-d_{\theta},\nu}$ is a random variable with the $F$ distribution on $n-d_{\theta}$ and $\nu$ degrees of freedom. It follows that the inverse lengths are of order $\|T^{(i)}\|^{-1}=O_{p}((n-d_{\theta})^{-1/2})$ uniformly in $i$. This allows control over the remainder terms. That $\bigl{\langle}T^{(i)}/\|T^{(i)}\|_{2},T^{(j)}/\|T^{(j)}\|_{2}\bigr{\rangle}$ has the same distribution as $\bigl{\langle}Z^{(i)}/\|Z^{(i)}\|_{2},Z^{(j)}/\|Z^{(j)}\|_{2}\bigr{\rangle}$ up to this order follows from the discussion of §3.2. It follows that
$\hat{R}_{n}\stackrel{{\scriptstyle d}}{{=}}R_{n}+O_{p}\Bigl{(}\max\Bigl{\{}k\nu^{-1/2}(n-d_{\theta})^{-1},k\nu^{-1}(n-d_{\theta})^{-1/2}\Bigr{\}}\Bigr{)}=R_{n}+O_{p}(kn^{-3/2}).$
Appendix B An alternative proof of Proposition 4.4
As noted in the proof of Proposition 4.4, the result as stated is a direct application of Polya’s theorem. An alternative derivation is via an application of the following proposition, which replaces the $o(1)$ term in the statement of Proposition 4.4. These rates may well be very conservative. On choosing $\zeta=\gamma^{2}/8$ in proposition B.1 and noting that
$\frac{b^{c^{2}/2\gamma+c\sqrt{\zeta/\gamma}-1}}{\sqrt{\log b}}\rightarrow 0\text{, and }\frac{2}{b^{(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}/4}}\rightarrow 0$
for any $c<\sqrt{\gamma/2}$ we see that the error term in Proposition B.1 tends to zero, recovering Proposition 4.4.
Proposition B.1.
Let $\hat{\sigma}^{2}$ be an estimator, independent of $R^{2}$, such that with $b_{n}=\gamma(n-d_{\theta})$ for some $0<\gamma\leq 1$, and $\epsilon_{n}=c\sqrt{\log(b_{n})/b_{n}}$ for some $c<2\sqrt{\zeta/\gamma}$ and $\zeta>0$,
$\mathop{\mathrm{pr}}\Bigl{(}\Bigl{|}\frac{\hat{\sigma}^{2}}{\sigma^{2}}-1\Bigr{|}\leq\epsilon_{n}\Bigr{)}\geq 1-r_{n}(\epsilon_{n}).$
Then,
$\sup_{v>0}|\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)|\leq 3r_{n}(\epsilon_{n})+O\left(\frac{b^{c^{2}/2+c\sqrt{\zeta}-1}}{\sqrt{\log b}}\right)+O\left(\frac{2}{b^{(2\sqrt{\zeta}-c)^{2}/4}}\right).$
Proof.
Here we outline the argument. Let $\epsilon_{n}=c\sqrt{\log(\gamma b)/\gamma b}$ with $c<2\sqrt{\zeta/\gamma}$ and consider
$\sup_{v\in\mathbb{R}}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$ $\leq\sup_{|v-b|<2\sqrt{b\zeta\log(b)}+2\zeta\log(b)}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$
$+\sup_{|v-b|\geq 2\sqrt{b\zeta\log(b)}+2\zeta\log(b)}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}.$
We use a Riemann approximation on the bulk of the distribution for values of $v$ close to $b$, and control the truncation error for very large values of $v$, as the Riemann approximation becomes harder to control for large values of $v$. Lemmas B.1 and B.2 give
$\sup_{v\in\mathbb{R}}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}\leq 3r_{n}(\epsilon_{n})+O\left(\frac{b^{c^{2}/2\gamma+c\sqrt{\zeta/\gamma}-1}}{\sqrt{\log b}}\right)+\frac{2}{b^{(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}/4}},$
as, in the notation defined there, $\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))\leq\mathop{\mathrm{pr}}(\mathcal{E}^{c}(\epsilon_{n}))=r_{n}(\epsilon_{n})$. ∎
Lemma B.1.
For $10<\epsilon_{n}<$ 1, introduce the events
$\mathcal{E}_{+}(\epsilon_{n}):\{\hat{\sigma}^{2}/\sigma^{2}\leq 1+\epsilon_{n}\},\quad\quad\mathcal{E}(\epsilon_{n}):=\{|\hat{\sigma}^{2}/\sigma^{2}-1|\leq\epsilon_{n}\}.$
Let $b=n-d_{\theta}$, the other notation being that of Proposition 4.4. Then for $\epsilon_{n}=c\sqrt{\log(\gamma b)/\gamma b}$, for some $c<2\sqrt{\zeta/\gamma}$ and $0<\gamma\leq 1$,
$\sup_{|v-b|<2\sqrt{\zeta b\log(b)}+2\zeta\log(b)}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$
$\leq\mathop{\mathrm{pr}}(\mathcal{E}^{c}(\epsilon_{n}))+2\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))+O\biggl{(}\frac{b^{c^{2}/2\gamma+c\sqrt{\zeta/\gamma}-1}}{\sqrt{\log b}}\biggr{)}.$
Proof.
On the event $\mathcal{E}_{+}(\epsilon_{n})$, $\{\hat{V}\leq v\}$ implies $\{V\leq v(1+\epsilon_{n})\}$, giving the probability bound
$\mathop{\mathrm{pr}}(\hat{V}\leq v)\leq\mathop{\mathrm{pr}}(V\leq v(1+\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}_{+}(\epsilon_{n}))+\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))$ (B.1)
For a lower bound, write
$\hat{V}=V(\sigma^{2}/\hat{\sigma}^{2}-1)+V\leq V(1+|\sigma^{2}/\hat{\sigma}^{2}-1|).$
Since $\mathcal{E}(\epsilon_{n})$ can be written as
$\mathcal{E}(\epsilon_{n})=\Bigl{\{}\frac{\epsilon_{n}}{1+\epsilon_{n}}\leq\frac{\sigma^{2}}{\hat{\sigma}^{2}}-1\leq\frac{\epsilon_{n}}{1-\epsilon_{n}}\Bigr{\}},$
on the event $\mathcal{E}(\epsilon_{n})$,
$\{\hat{V}\leq v\}\supseteq\Bigl{\{}V\leq\frac{v}{1+\epsilon_{n}/(1-\epsilon_{n})}\Bigr{\}}=\{V\leq v(1-\epsilon_{n})\}.$
This implies the lower bound
$\mathop{\mathrm{pr}}(\hat{V}\leq v)\geq\mathop{\mathrm{pr}}(V\leq v(1-\epsilon_{n}))\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))$ (B.2)
By combining (B.1) and (B.2)
$\sup_{|v-b|<2\sqrt{\zeta b\log(b)}+2\zeta\log(b)}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$
$\leq\sup_{|v-b|<2\sqrt{\zeta b\log(b)}+2\zeta\log(b)}\sup_{\bar{v}:|v-\bar{v}|\leq v\epsilon_{n}}U(v,\bar{v},\epsilon_{n}),$
where
$U(v,\bar{v},\epsilon_{n})$ $=$ $\max\Bigl{\{}\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(\epsilon_{n}))-\mathop{\mathrm{pr}}(V\leq v)\bigr{|},$
$\quad\quad\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}_{+}(\epsilon_{n}))-\mathop{\mathrm{pr}}(V\leq v)\Bigr{|}+\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n}))\bigr{\}}.$
The two terms in the maximum can be upper bounded as
$|\displaystyle\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}(t))-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}\leq$$\;\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}+\mathop{\mathrm{pr}}(\mathcal{E}^{c}(\epsilon_{n})),$
$\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})\,\mathop{\mathrm{pr}}(\mathcal{E}_{+}(t))-\mathop{\mathrm{pr}}(V\leq v)\Bigr{|}+\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(t))\leq$ $\;\bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\Bigr{|}+2\mathop{\mathrm{pr}}(\mathcal{E}_{+}^{c}(\epsilon_{n})).$
We thus need an upper bound for $|\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)|$. Note that
$\mathop{\mathrm{pr}}(V\leq\bar{v})=\mathop{\mathrm{pr}}\left(V\leq Tv\right),$
for some value of $T$ in $(1-\epsilon_{n},1+\epsilon_{n})$, and $V$ is a chi-square random variable with $b$ degrees of freedom. We consider four cases:
Case 1: $T\geq 1$, $v\geq b$. Since the mode of the density of $V$ is at $b-2$, we use the following Riemann approximation to the relevant intractable integral
$\mathop{\mathrm{pr}}\left(V\leq Tv\right)-\mathop{\mathrm{pr}}(V\leq v)$
[…]
$\leq\frac{1}{2^{b/2}\Gamma(b/2)}\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}(v+i)^{b/2-1}\exp\left(-\frac{v}{2}-\frac{i}{2}\right)$
$=\frac{v^{b/2-1}}{2^{b}\Gamma(b/2)}\exp\left(-\frac{v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i}{v}\right)^{b/2-1}\exp\left(-\frac{i}{2}\right)$
$\leq\frac{v^{b/2-1}}{2^{b-1}\Gamma(b/2)}\exp\left(-\frac{v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i}{v}\right)^{b/2}\exp\left(-\frac{i}{2}\right),$
We bound the sum by
$\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\biggl{(}1+\frac{i}{v}\biggr{)}^{b/2}\exp\biggl{(}-\frac{i}{2}\biggr{)}$
$\leq 2\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\biggl{\{}\exp\biggl{(}\frac{1}{2}-\frac{1}{2}\biggr{)}\biggr{\}}^{i}\leq 2\lceil\epsilon_{n}v\rceil$
as $(1+i/v)^{b}\leq\exp(i)$ and $v\geq b$. It remains to upper bound the multiplicative term
$2\lceil\epsilon_{n}v\rceil\frac{v^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\left(-\frac{v}{2}\right)$
$\leq\frac{4\lceil\epsilon_{n}v\rceil}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{v}{2}+\Bigl{(}\frac{b}{2}-1\Bigr{)}\log v-\frac{b}{2}\log b+\frac{b}{2}\biggr{\}}$
$\leq\frac{4\lceil\epsilon_{n}v\rceil}{\sqrt{\pi b}}\exp\Bigl{\{}-\frac{v}{2}+\frac{b}{2}(\log(v/b)+1)-\log v\Bigr{\}}$
$\leq\frac{4\lceil\epsilon_{n}v\rceil}{\sqrt{\pi b}}\exp\left(-\log v\right)$
$\leq\frac{\lceil\epsilon v\rceil}{v\sqrt{\pi b}}\leq\frac{\epsilon_{n}}{\sqrt{\pi b}}+\frac{1}{\sqrt{\pi}b^{3/2}}=O\biggl{(}\frac{\sqrt{\log b}}{b^{3/2}}\biggr{)}.$
Case 2: $T<1$, $v\geq b$. By the log-convexity of the gamma distribution for all $x>0$ and $t\in[0,1]$ the gamma density function $\gamma$ satisfies
$\gamma(x+t)\leq\frac{1}{2^{b/2}\Gamma(b/2)}(x+1)^{b/2-1}\exp(-x/2).$
Based on this estimate the Riemann approximation to the relevant integral is
$\mathop{\mathrm{pr}}\left(V\leq v\right)-\mathop{\mathrm{pr}}(V\leq Tv)$
[…]
$\leq\frac{1}{2^{b/2}\Gamma(b/2)}\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}((1-\epsilon_{n})v+i+1)^{b/2-1}\exp\left(-\frac{(1-\epsilon_{n})v}{2}-\frac{i}{2}\right)$
=$=\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2-1}\exp\left(-\frac{i}{2}\right)$
$\leq\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\left(-\frac{(1-\epsilon_{n})v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2}\exp\left(-\frac{i}{2}\right).$ (B.3)
In this expression,
$\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2}$ $=\left(1+\frac{i+1}{b}\frac{b}{(1-\epsilon_{n})v}\right)^{b/2}$
$\leq\exp\left(\frac{ib}{2(1-\epsilon_{n})v}+\frac{b}{2(1-\epsilon_{n})v}\right)$
$\leq e^{1/4}\exp\left(\frac{ib(1+\epsilon_{n})}{2(1+\epsilon_{n}^{2})v}\right\}$
$\leq e^{1/4}\exp\left(\frac{ib(1+\epsilon_{n})}{2v}\right\},$
where the last inequality is because $v\geq b$. Thus
$\left(1+\frac{i+1}{(1-\epsilon_{n})v}\right)^{b/2}\exp\left(-\frac{i}{2}\right)\leq e^{1/4}\exp(\epsilon_{n}i/2)$
in equation (B) and the partial geometric series is bounded, for large enough $b$, by
$\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\biggl{(}1+\frac{i}{(1-\epsilon_{n})v}\biggr{)}^{b/2}\exp\biggl{(}-\frac{i}{2}\biggr{)}\leq\frac{\exp(\frac{\lceil\epsilon_{n}v\rceil}{2}\epsilon_{n})-1}{\exp(\frac{1}{2}\epsilon_{n})-1}\leq 2e^{1/2}\epsilon_{n}^{-1}\exp\left(\frac{c^{2}\log(\gamma b)}{2\gamma}\right).$
This follows because $v\leq b+2\sqrt{\zeta b\log b}+2\zeta\log(b)$ and $v/b-1\rightarrow 0$.
Consider the multiplicative term in (B). For $b$ sufficiently large,
$\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\biggl{\{}-\frac{(1-\epsilon_{n})v}{2}\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{(1-\epsilon_{n})v}{2}+\Bigl{(}\frac{b}{2}-1\Bigr{)}\log((1-\epsilon_{n})v)-\frac{b}{2}\log(b)+\frac{b}{2}-\frac{1}{12b+1}\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}\left(\frac{b}{2}\right)\log\left\{\frac{(1-\epsilon_{n})v)}{b}\right\}+\frac{b}{2}-\frac{(1-\epsilon_{n})v}{2}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}\frac{b}{2}\Bigl{(}\log(1-\epsilon_{n})+\log(v/b)+1-\frac{(1-\epsilon_{n})v}{b}\Bigr{)}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{\epsilon_{n}b}{2}(1-\frac{v}{b})-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2\exp(\epsilon_{n}\sqrt{\zeta b\log(b)}+\epsilon_{n}\zeta\log(b))}{\sqrt{\pi b}(1-\epsilon_{n})b}$
$\leq\frac{4\exp(c\sqrt{\zeta/\gamma}\log(b))}{\sqrt{\pi}b^{3/2}},$
where we have used the lower bound for Stirling’s approximation. Thus the overall error in Case 2 is
$O\biggl{(}\frac{b^{c^{2}/2\gamma+c\sqrt{\zeta/\gamma}-1}}{\sqrt{\log b}}\biggr{)}.$
Case 3: $T\geq 1$, $v<b$. Similar arguments give
$\mathop{\mathrm{pr}}\left(V\leq Tv\right)-\mathop{\mathrm{pr}}(V\leq v)$ $\leq\frac{v^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\left(-\frac{v}{2}\right)\sum^{\lceil\epsilon_{n}v\rceil}_{i=1}\left(1+\frac{i+1}{v}\right)^{b/2}\exp\left(-\frac{i}{2}\right),$
and, for sufficiently large $b$,
$\left(1+\frac{i+1}{v}\right)^{b/2}\exp\left(-\frac{i}{2}\right)\leq e^{1/4}\exp\left\{\{\sqrt{\zeta\log(b)}b^{-1/2}\}i/2\right\},$
so that
$\sum^{\lceil\epsilon_{n}v\rceil}_{i=1}\left(1+\frac{i}{(1-\epsilon_{n})v}\right)^{b/2}\exp\left(-\frac{i}{2}\right)$$\leq$ $2\exp\biggl{(}\frac{\lceil\epsilon_{n}v\rceil}{2}b^{-1/2}\sqrt{\zeta\log b}\biggr{)}$
$\leq$ $4\biggl{(}\frac{b}{\zeta\log b}\biggr{)}^{1/2}\exp\biggl{(}\frac{c\sqrt{\zeta}\log b}{2\sqrt{\gamma}}\biggr{)}.$
The multiplicative term satisfies
$\frac{\{v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\biggl{(}-\frac{v}{2}\biggr{)}$$\leq$ $\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{\epsilon_{n}b}{2}\Bigl{(}1-\frac{v}{b}\Bigr{)}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq$ $\frac{2}{\sqrt{\pi b}}\exp\left(-\log((1-\epsilon_{n})v)\right)$
$=$ $\frac{2}{\sqrt{\pi b}(1-\epsilon_{n})(b-\sqrt{2}{\zeta\log b}-2\zeta\log b)}.$
Therefore the error for this case is
$O\biggl{(}\frac{b^{c\sqrt{\zeta}/2\sqrt{\gamma}-1}}{\sqrt{\log b}}\biggr{)}.$
Case 4: $T<1$, $v<b$. In this case
$\mathop{\mathrm{pr}}(V\leq v)-\mathop{\mathrm{pr}}(V\leq Tv)$
$\leq\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\biggl{(}-\frac{(1-\epsilon_{n})v}{2}\biggr{)}\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\biggl{(}1+\frac{i}{(1-\epsilon_{n})v}\biggr{)}^{b/2}\exp\biggl{(}-\frac{i}{2}\biggr{)},$
and
$\biggl{(}1+\frac{i}{(1-\epsilon_{n})v}\biggr{)}^{b/2}\exp\left(-\frac{i}{2}\right)\leq e^{1/2}\exp\biggl{\{}\frac{i}{2}(\epsilon_{n}+\{\sqrt{\zeta\log(b)}b^{-1/2}\}+\epsilon_{n}\{\sqrt{\zeta\log(b)}b^{-1/2}\})\biggr{\}}.$
Therefore, for $b$ sufficiently large,
$\sum^{\lceil\epsilon_{n}v\rceil}_{i=0}\biggl{(}1+\frac{i}{(1-\epsilon_{n})v}\biggr{)}^{b/2}\exp\left(-\frac{i}{2}\right)$
$\leq 2\frac{\exp\left\{\frac{\lceil\epsilon_{n}v\rceil}{2}(\epsilon_{n}+\{\sqrt{\zeta\log(b)}b^{-1/2}\}+\epsilon_{n}\{\sqrt{\zeta\log(b)}b^{-1/2}\}\right\}-1}{\exp\left\{\frac{1}{2}(\epsilon_{n}+\{\sqrt{\zeta\log(b)}b^{-1/2}\}+\epsilon_{n}\{\sqrt{\zeta\log(b)}b^{-1/2}\}\right\}-1}$
$\leq\frac{4e\sqrt{b}}{(c/\sqrt{\gamma}+\sqrt{\zeta})\log(\gamma b)}\exp\biggl{(}\frac{c^{2}/\gamma+c\sqrt{\zeta/\gamma}}{2}\log b\biggr{)}.$
Since $v\geq b-2\sqrt{\zeta b\log(b)}-2\zeta\log(b)$, the multiplicative term satisfies
$\frac{\{(1-\epsilon_{n})v\}^{b/2-1}}{2^{b/2-1}\Gamma(b/2)}\exp\biggl{\{}-\frac{(1-\epsilon_{n})v}{2}\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{(1-\epsilon_{n})v}{2}+\Bigl{(}\frac{b}{2}-1\Bigr{)}\log((1-\epsilon_{n})v)-\frac{b}{2}\log(b)+\frac{b}{2}-\frac{1}{12b+1}\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}\Bigl{(}\frac{b}{2}\Bigr{)}\log\Bigl{(}\frac{(1-\epsilon_{n})v}{b}\Bigr{)}+\frac{b}{2}-\frac{(1-\epsilon_{n})v}{2}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}\frac{b}{2}\Bigl{(}\log(1-\epsilon_{n})+\log(v/b)+1-\frac{(1-\epsilon_{n})v}{b}\Bigr{)}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}}\exp\biggl{\{}-\frac{\epsilon_{n}b}{2}\Bigl{(}1-\frac{v}{b}\Bigr{)}-\log((1-\epsilon_{n})v)\biggr{\}}$
$\leq\frac{2}{\sqrt{\pi b}(1-\epsilon_{n})\{b-2\sqrt{\zeta b\log b}-2\zeta\log(b)\}}=O\Bigl{(}\frac{1}{b^{3/2}}\Bigr{)},$
giving an error of
$O\biggl{(}\frac{b^{\frac{c^{2}/\gamma+c\sqrt{\zeta/\gamma}}{2}-1}}{\sqrt{\log b}}\biggr{)},$
which is dominated by the error in Case 2. ∎
The following bounds are used to control the truncation error in Lemma B.2.
Proposition B.2 (Laurant and Massart 2000, Lemma 1).
For a random variable $Z$ distributed as a Chi-square with $b$ degrees of freedom the following tail bounds hold:
$\mathop{\mathrm{pr}}(Z\geq b+2\sqrt{bt}+2t)\leq\exp(-t)$
$\mathop{\mathrm{pr}}(Z\leq b-2\sqrt{bt})\leq\exp(-t).$
Lemma B.2.
Let $\epsilon_{n}=c\sqrt{\log(\gamma b)/\gamma b}$ and $c<2\sqrt{\zeta/\gamma}$, then
$\sup_{|v-b|\geq 2\sqrt{b\zeta\log(b)}+2\zeta\log(b)}\bigl{|}\mathop{\mathrm{pr}}(\hat{V}\leq v)-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}\leq\frac{2}{b^{(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}/4}}.$
Proof.
$\sup_{v\geq b+2\sqrt{\zeta b\log b}+2\zeta\log b}\ \bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$
$\leq\sup_{v\geq b+2\sqrt{\zeta b\log b}+2\zeta\log b}\ \mathop{\mathrm{pr}}(V>(1-\epsilon_{n})v)+\sup_{v\geq b+2\sqrt{\zeta b\log b}+2\zeta\log b}\ \mathop{\mathrm{pr}}(V>v)$
$\leq 2\mathop{\mathrm{pr}}\Bigl{(}V>(1-\epsilon_{n})\bigl{\{}b+2\sqrt{\zeta b\log b}+2\zeta\log b\bigr{\}}\Bigr{)}$
$=2\mathop{\mathrm{pr}}\left(V>b+(2\sqrt{\zeta}-c/\sqrt{\gamma})\sqrt{b\log b}+(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}\log b\right)\leq\frac{2}{b^{(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}/4}},$
since $\epsilon_{n}b=c(b\log(b)/\gamma)^{1/2}$ and $\epsilon_{n}\{2\sqrt{\zeta b\log b}+2\zeta\log b\}<c\log(b)/\gamma$ eventually for sufficiently large $b$. By a similar argument for the left tail we have
$\sup_{v\leq b-2\sqrt{\zeta b\log b}-2\zeta\log b}\ \bigl{|}\mathop{\mathrm{pr}}(V\leq\bar{v})-\mathop{\mathrm{pr}}(V\leq v)\bigr{|}$
$\leq\sup_{v\leq b-2\sqrt{\zeta b\log b}-2\zeta\log b}\ \mathop{\mathrm{pr}}(V<(1+\epsilon_{n})v)+\sup_{v\leq b-2\sqrt{\zeta b\log b}-2\zeta\log b}\ \mathop{\mathrm{pr}}(V<v)$
$\leq 2\mathop{\mathrm{pr}}\Bigl{(}(V<(1+\epsilon_{n})\{b-2\sqrt{\zeta b\log b}-2\zeta\log b\}\Bigr{)}$
$=2\mathop{\mathrm{pr}}\Bigl{(}V<b-(2\sqrt{\zeta}-c/\sqrt{\gamma})\sqrt{b\log b}\Bigr{)}\leq\frac{2}{b^{(2\sqrt{\zeta}-c/\sqrt{\gamma})^{2}/4}}.$